### v1

# Memory-augmented Transformers can implement Linear First-Order Optimization Methods

Abstract
We show that memory-augmented Transformers (Memformers) can implement linear first-order optimization methods such as conjugate gradient descent, momentum methods, and more generally, methods that linearly combine past gradients. Building on prior work that demonstrates how Transformers can simulate preconditioned gradient descent, we provide theoretical and empirical evidence that Memformers can learn more advanced optimization algorithms. Specifically, we analyze how memory registers in Memformers store suitable intermediate attention values allowing them to implement algorithms such as conjugate gradient. Our results show that Memformers can efficiently learn these methods by training on random linear regression tasks, even learning methods that outperform conjugate gradient. This work extends our knowledge about the algorithmic capabilities of Transformers, showing how they can learn complex optimization methods.
1 Introduction
In-context learning (ICL) allows large language models (LLMs) to generate contextually appropriate outputs based solely on examples and queries provided in a prompt, without requiring any parameter adjustments (Brown, 2020; Liu et al., 2021; Lu et al., 2021; Wei et al., 2022; Wu et al., 2022). This remarkable ability has spurred research into understanding how Transformers can implement algorithms (Brown, 2020; Achiam et al., 2023; Touvron et al., 2023), with recent studies focusing on their capability to simulate optimization algorithms (Dai et al., 2022; Von Oswald et al., 2023a; Garg et al., 2022; Akyürek et al., 2022). Transformers have been shown to implement gradient-based optimization during their forward pass, such as preconditioned gradient descent for linear regression tasks (Dai et al., 2022; Mahankali et al., 2023; Ahn et al., 2024).
More recently, studies have demonstrated that Transformers can learn even more advanced optimization methods. For instance, Fu et al. (2023) showed that Transformers exhibit convergence rates comparable to Iterative Newton’s Method, a higher-order optimization technique that converges exponentially faster than gradient descent for in-context linear regression. Additionally, Vladymyrov et al. (2024) proved that Transformers can, in fact, learn a variant of gradient descent that approximates second-order methods, such as $\mathrm{GD}^{++}$, achieving convergence rates similar to Newton’s method. These findings lead to the central question of our paper:
Can Transformers efficiently learn more advanced gradient-based optimization methods?
We hope that by addressing this question, we can reveal some of the representational power of Transformers as “algorithm learners,” and thus further motivate the use of machine learning for discovering new optimization algorithms. To make our question more precise, we focus on learning the entire class of gradient-based algorithms obtained by linearly combining past gradients, known as Linear First-Order Methods (LFOMs) (Goh, 2017), where the $k+1^{\text{st}}$ iterate is
$w^{k+1}=w^{0}+\sum_{i=0}^{k}\Gamma_{i}^{k}\nabla f(w^{i}),$ (1.1)
and where $\{\Gamma_{i}^{k}\}_{i=0}^{k}$ are diagonal matrices. Model (1.1) is quite general, as it includes, as special cases, standard methods such as gradient descent (GD), momentum GD, Nesterov’s accelerated gradient, conjugate gradient, and in a stochastic setting, AdaGrad, ADAM, among others.
Our key insight for efficiently learning LFOMs is to leverage memory-augmented Transformers, known as Memformers (Wu et al., 2020; Xu et al., 2021), which retain intermediate attention values across layers. This memory can enable them to store past gradients, facilitating the execution of advanced first-order methods such as conjugate gradient descent and momentum methods. The same mechanism allows Memformers to implement more general LFOMs.
While unconditional learning of gradient methods remains out of reach, we build on related work demonstrating that Transformers can learn gradient descent in the context of linear regression tasks (Garg et al., 2022; Akyürek et al., 2022; Von Oswald et al., 2023a; Ahn et al., 2024; Zhang et al., 2024). Inspired by these findings, and extending the work of (Ahn et al., 2024), we conduct a theoretical analysis of the loss landscape for memory-augmented linear Transformers that omit softmax activation (Schlag et al., 2021; Von Oswald et al., 2023a; Ahn et al., 2024).
In summary, our main contributions are as follows:
Main Contributions
  1. (A)
    Theoretical and Experimental Evidence. We provide theoretical justification and empirical results showing that Memformers trained on linear regression tasks learn LFOMs (and even quasi-Newton methods like $\mathrm{GD}^{++}$ (Von Oswald et al., 2023a) for novel in-context data.
  2. (B)
    Performance Beyond Conjugate Gradient Descent. We demonstrate that LFOMs learned by Memformers significantly outperform conjugate gradient descent on training data, while remaining competitive on test data, indicating good generalization performance.
  3. (C)
    Multi-Headed Attention Improves Performance. We show empirically that multi-headed attention improves Memformers’ test performance, and offer theoretical justification for why increasing attention heads enhances loss performance on test data.
The main objective of this paper is to investigate the potential of augmented Transformers to learn optimization algorithms. We are not advocating for Transformers as substitutes for established optimization methods, but rather, we aim to kindle further exploration into their algorithmic capabilities. We believe our results contribute to understanding how augmented Transformers could help with optimization, which may ultimately lead to new and practical gradient-based algorithms.
1.1 Related Work
Research on Transformers is extremely active, and we cannot hope to fully capture the breadth of the related literature. Below, we summarize the most immediately relevant topics.
In-Context Learning. The ability of Transformer models to perform in-context learning (ICL) has been extensively studied since its introduction by Brown (2020). Subsequent works have explored how these models adapt to new tasks without requiring parameter updates (Xie et al., 2021; Von Oswald et al., 2023b; Hahn and Goyal, 2023; Liu et al., 2021; Lu et al., 2021; Wei et al., 2022; Wu et al., 2022). This foundational research has paved the way for studies investigating how Transformers can implement specific algorithms, such as gradient-based methods.
Gradient-Based Methods in Transformers. Garg et al. (2022) analyze the learning of gradient descent within Transformers, particularly in the context of ICL for linear functions. Empirical studies (Garg et al., 2022; Akyürek et al., 2022; Von Oswald et al., 2023a) have shown that Transformers can learn gradient descent after being trained on random linear regression tasks. Expanding on these results, Von Oswald et al. (2023a); Ahn et al. (2024) demonstrate that Transformers can implement preconditioned gradient descent for solving linear regression problems presented in input prompts. Notably, these works—as well as ours—utilize Linear Transformers as discussed in (Schlag et al., 2021; Von Oswald et al., 2023a; Ahn et al., 2023).
Higher-Order Optimization Methods in Transformers. Transformers have also been shown to learn higher-order optimization techniques, such as Newton’s method, expanding their capabilities beyond first-order methods (Fu et al., 2023; Giannou et al., 2024; Vladymyrov et al., 2024).
Memory-Augmented Transformers (Memformers). Memformers were introduced by Wu et al. (2020); Xu et al. (2021). These models retain intermediate attention values across layers through memory registers, enabling more complex computations and optimization methods to be learned. While significant progress has been made in understanding how Transformers can learn gradient descent, their potential for implementing more sophisticated LFOMs remains largely unexplored. Our work addresses this gap by showing how Memformers can efficiently implement a wide range of advanced first-order and quasi-second-order optimization techniques, including CGD and momentum methods, thereby pushing the boundaries of Transformer-based architectures.
2 Background and Problem Setup
2.1 Linear Transformers on Random Linear Regression
We follow the setup of training Transformers on random instances of linear regression, following the prior works (Garg et al., 2022; Akyürek et al., 2022; Von Oswald et al., 2023a; Ahn et al., 2024). We largely use the notation and formal setup of (Ahn et al., 2024), which we now proceed to recall.
Data Distribution. Let $\mathbf{x}(i)\in\mathbb{R}^{d}$ represent covariates drawn independently from a distribution $\mathcal{D}_{\mathbf{X}}$, and let $\mathbf{w}^{*}\in\mathbb{R}^{d}$ be drawn from $\mathcal{D}_{\mathbf{W}}$. The matrix of covariates $\mathbf{X}\in\mathbb{R}^{(n+1)\times d}$ contains rows $\mathbf{x}(i)$. The responses are $\mathbf{y}=[\langle\mathbf{x}(1),\mathbf{w}^{*}\rangle,\dots,\langle\mathbf{x}(n),\mathbf{w}^{*}\rangle]\in\mathbb{R}^{n}$. Define the input matrix $\mathbf{Z}_{0}$ as:
$\mathbf{Z}_{0}=\begin{bmatrix}\mathbf{x}(1)&\mathbf{x}(2)&\cdots&\mathbf{x}(n)&\mathbf{x}(n+1)\\ \mathbf{y}(1)&\mathbf{y}(2)&\cdots&\mathbf{y}(n)&0\end{bmatrix}\in\mathbb{R}^{(d+1)\times(n+1)},$ (2.1)
where the zero corresponds to the unknown response for $\mathbf{x}(n+1)$. The task is to predict $(\mathbf{w}^{*})^{\top}\mathbf{x}(n+1)$ using $\mathbf{Z}_{0}$. The training data consists of pairs $(\mathbf{Z}_{0},(\mathbf{w}^{*})^{\top}\mathbf{x}(n+1))$ for $\mathbf{x}(i)\sim\mathcal{D}_{\mathbf{X}}$ and $\mathbf{w}^{*}\sim\mathcal{D}_{\mathbf{W}}$.
Self-Attention Without Softmax. We focus on the linear self-attention layer, building on (Schlag et al., 2021; Von Oswald et al., 2023a). Let $\mathbf{Z}\in\mathbb{R}^{(d+1)\times(n+1)}$ be the input matrix of $n+1$ tokens in $\mathbb{R}^{d+1}$. Standard self-attention layer is defined as
$\text{Attn}_{\text{smax}}(\mathbf{Z}):=W_{v}\mathbf{Z}M\cdot\text{smax}(\mathbf{Z}^{\top}W_{k}^{\top}W_{q}\mathbf{Z}),$ (2.2)
where $W_{v},W_{k},W_{q}\in\mathbb{R}^{(d+1)\times(d+1)}$ are weight matrices, and $\text{smax}(\cdot)$ denotes the column-wise softmax. The masking matrix $M$ ensures that the label for $\mathbf{x}(n+1)$ is excluded is given by
$M=\begin{bmatrix}\mathbf{I}_{n}&0\\ 0&0\end{bmatrix}\in\mathbb{R}^{(n+1)\times(n+1)}.$ (2.3)
Omitting softmax, the attention mechanism becomes
$\text{Attn}_{P,Q}(\mathbf{Z}):=P\mathbf{Z}M(\mathbf{Z}^{\top}Q\mathbf{Z}),$ (2.4)
where $P=W_{v}$ and $Q=W_{k}^{\top}W_{q}$. This simplified form, as shown in Ahn et al. (2024), can implement preconditioned gradient descent, and it is the one we also use.
Transformer Architecture. As in the related work, we also simplify the Transformer to consider only attention layers, using $L$ layers of linear self-attention with a residual connection. Therefore, for each layer $\ell$, the output is updated as
$\mathbf{Z}_{\ell+1}=\mathbf{Z}_{\ell}+\frac{1}{n}\text{Attn}_{P_{\ell},Q_{\ell}}(\mathbf{Z}_{\ell}),\quad\ell=0,1,\dots,L-1.$ (2.5)
Using updates (2.5), with the input $\mathbf{Z}_{0}$, the final transformer output is
$\text{TF}_{L}(\mathbf{Z}_{0};\{P_{\ell},Q_{\ell}\}_{\ell=0}^{L-1})=-[\mathbf{Z}_{L}]_{(d+1),(n+1)}.$ (2.6)
The set of parameters $\{P_{\ell},Q_{\ell}\}_{\ell=0}^{L-1}$ is then learned by minimizing the following training objective:
$f\left(\{P_{\ell},Q_{\ell}\}_{\ell=0}^{L-1}\right)=\mathbb{E}_{(\mathbf{Z}_{0},\mathbf{w}^{*})}\left[\left(\text{TF}_{L}(\mathbf{Z}_{0})+(\mathbf{w}^{*})^{\top}\mathbf{x}(n+1)\right)^{2}\right].$ (2.7)
Here, the scaling factor $\frac{1}{n}$ is used only for ease of notation and does not influence the expressive power of the Transformer.
We will utilize the following lemma from Ahn et al. (2024), which demonstrates that multi-layer Transformers simulate preconditioned gradient descent under suitable parameterization. We have provided the full proof of this Lemma 2.11 in the Appendix for completeness.
$P_{\ell}=\begin{bmatrix}\mathbf{B}_{\ell}=0_{d\times d}&0\\ 0&1\end{bmatrix},\quad Q_{\ell}=-\begin{bmatrix}\mathbf{A}_{\ell}&0\\ 0&0\end{bmatrix},\quad\mathbf{A}_{\ell},\mathbf{B}_{\ell}\in\mathbb{R}^{d\times d}.$ (2.8)
Lemma 1 (Lemma 1, Ahn et al. (2024)).
Consider an $L$-layer linear transformer parameterized by $\mathbf{A}_{0},\dots,\mathbf{A}_{L-1}$, as in (2.8). Let $y_{\ell}^{(n+1)}$ be the $(d+1,n+1)$-th entry of the $\ell$-th layer output, i.e., $y_{\ell}^{(n+1)}=[\mathbf{Z}_{\ell}]_{(d+1),(n+1)}$ for $\ell=1,\dots,L$.
$y_{\ell}^{(n+1)}=-\langle\mathbf{x}^{(n+1)},\mathbf{w}_{\ell}^{\mathrm{gd}}\rangle,$ (2.9)
where the sequence $\{\mathbf{w}_{\ell}^{\mathrm{gd}}\}$ is defined as $\mathbf{w}_{0}^{\mathrm{gd}}=0$ and for $\ell=1,\dots,L-1$:
$\mathbf{w}_{\ell+1}^{\mathrm{gd}}=\mathbf{w}_{\ell}^{\mathrm{gd}}-\mathbf{A}_{\ell}\nabla R_{\mathbf{w}^{*}}(\mathbf{w}_{\ell}^{\mathrm{gd}}),$ (2.10)
with the empirical least-squares loss (with $\mathbf{X}:=[\mathbf{x}^{(1)},\ldots,\mathbf{x}^{(n)}]\in\mathbb{R}^{d\times n}$):
$R_{\mathbf{w}^{*}}(\mathbf{w}):=\frac{1}{2n}\|\mathbf{X}^{\top}\mathbf{w}-\mathbf{X}^{\top}\mathbf{w}^{*}\|^{2}=\frac{1}{2n}(\mathbf{w}-\mathbf{w}^{*})^{\top}\mathbf{X}\mathbf{X}^{\top}(\mathbf{w}-\mathbf{w}^{*}).$ (2.11)
Note: The Hessian of $R_{\mathbf{w}^{*}}(\mathbf{w})$ is given by:
$\nabla^{2}R_{\mathbf{w}^{*}}(\mathbf{w})=\frac{1}{n}\mathbf{X}\mathbf{X}^{\top}.$ (2.12)
2.2 Linear First-Order Methods
Linear First-Order Methods (LFOMs) (Goh, 2017) are a class of optimization algorithms that lineary combine past gradients for minimizing smooth objective functions. They iteratively update a parameter vector $\mathbf{w}$ using the gradient of the objective function. The general update rule is
$\mathbf{w}^{k+1}=\mathbf{w}^{k}+\alpha_{k}\mathbf{d}^{k},$ (2.13)
where $\alpha_{k}$ is the step size and $\mathbf{d}^{k}$ is the update direction, typically related to the gradient $\nabla f(\mathbf{w}^{k})$. Algorithms within this family differ in how they compute $\mathbf{d}^{k}$ and choose $\alpha_{k}$.
LFOMs can be expressed in a cumulative form. For gradient descent, unrolling (2.13) we get
$\mathbf{w}^{k+1}=\mathbf{w}^{0}-\alpha\sum\nolimits_{i=0}^{k}\nabla f(\mathbf{w}^{i}),$ (2.14)
while common momentum methods need an additional term incorporating past gradients, yielding
$\mathbf{w}^{k+1}=\mathbf{w}^{0}+\sum\nolimits_{i=0}^{k}\gamma_{i}^{k}\nabla f(\mathbf{w}^{i}),$ (2.15)
where the coefficients $\gamma_{i}^{k}$ weight previous gradients. More advanced methods, or general LFOMs, use diagonal matrices $\Gamma_{i}^{k}$ to coordinate-wise scale each gradient component, i.e.,
$\mathbf{w}^{k+1}=\mathbf{w}^{0}+\sum\nolimits_{i=0}^{k}\Gamma_{i}^{k}\nabla f(\mathbf{w}^{i}).$ (2.16)
Momentum Methods and Conjugate Gradient Descent (CGD)
Momentum methods accelerate convergence by incorporating a momentum term, modifying the gradient to account for past updates and achieving faster convergence in relevant directions. Conjugate Gradient Descent (CGD), on the other hand, is a first-order method optimized for quadratic minimization, serving as a benchmark for large-scale, sparse linear systems. After an initial steepest descent, CGD generates directions conjugate to previous ones, leading to faster convergence than standard gradient descent. Both are core methods within the LFOM class, summarized below:
Momentum Methods
1:Initialize $\mathbf{w}_{0}$, $\mathbf{v}_{0}=0$
2:for $n=1,2,\dots$ do
3:     Compute the gradient:
$\nabla f(\mathbf{w}_{n})$
4:     Update the velocity:
$\mathbf{v}_{n}=\beta\mathbf{v}_{n-1}-\eta\nabla f(\mathbf{w}_{n})$
5:     Update the iterate:
$\mathbf{w}_{n+1}=\mathbf{w}_{n}+\mathbf{v}_{n}$
6:end for
7:$\beta$: Momentum coefficient (controls the influence of past gradients)
8:$\eta$: Learning rate (scales the gradient step size)
Conjugate Gradient Descent (CGD)
1:Initialize $\mathbf{w}_{0}$, $\mathbf{s}_{0}=-\nabla f(\mathbf{w}_{0})$
2:for $n=1,2,\dots$ do
3:     Compute the steepest descent direction:
$\Delta\mathbf{w}_{n}=-\nabla f(\mathbf{w}_{n})$
4:     Compute the conjugacy coefficient:
$\gamma_{n}=\frac{\|\nabla f(\mathbf{w}_{n})\|^{2}}{\|\nabla f(\mathbf{w}_{n-1})\|^{2}}$
5:     Update the search direction:
$\mathbf{s}_{n}=\Delta\mathbf{w}_{n}+\gamma_{n}\mathbf{s}_{n-1}$
6:     Perform a line search:
$\alpha_{n}=\arg\min_{\alpha}f(\mathbf{w}_{n}+\alpha\mathbf{s}_{n})$
7:     Update the iterate:
$\mathbf{w}_{n+1}=\mathbf{w}_{n}+\alpha_{n}\mathbf{s}_{n}$
8:end for
Momentum methods provide fast convergence by accumulating gradient history and are widely used in modern optimization. CGD converges in at most $N$ iterations for quadratic functions, where $N$ is the number of variables, and is effective for ill-conditioned problems.
3 Memformers Implement LFOMs In-Context
Building on Lemma 2.11, we demonstrate how memory-augmented Transformers can implement Linear First-Order Methods (LFOMs) like Conjugate Gradient Descent (CGD) in-context. As noted in Ahn et al. (2024) (Subsection C.1), the term $\text{Attn}_{P_{\ell},Q_{\ell}}(\mathbf{Z}_{\ell})$ in the update for $\mathbf{Z}_{\ell+1}$ corresponds to the preconditioned gradient $\mathbf{A}_{\ell}\nabla R_{\mathbf{w}^{*}}(\mathbf{w}_{\ell}^{\mathrm{gd}})$ of the in-context loss (2.11).
3.1 Dynamic Memory for CGD-like Algorithms
With attention simulating preconditioned gradient descent, we now explore how memory-augmented Transformers can simulate Conjugate Gradient Descent (CGD). The dynamic memory mechanism enables recursive refinement of search directions, closely mimicking CGD’s iterative process.
Proposition 1.
A memory-augmented Transformer can implement Conjugate Gradient Descent (CGD) through a dynamic memory mechanism that recursively refines search directions, where the update rules are:
$\mathbf{R}_{\ell}$ $=\mathrm{Attn}_{P_{\ell},Q_{\ell}}(\mathbf{Z}_{\ell})+\gamma_{\ell}\mathbf{R}_{\ell-1},$ (3.1)
$\mathbf{Z}_{\ell+1}$ $=\mathbf{Z}_{\ell}+\alpha_{\ell}\frac{1}{n}\mathbf{R}_{\ell},$ (3.2)
where $\gamma_{\ell}$ and $\alpha_{\ell}$ control past update influence and step size
Proof Sketch.  CGD refines search directions using current gradients and previous directions. The Transformer simulates this by using $\text{Attn}_{P_{\ell},Q_{\ell}}(\mathbf{Z}_{\ell})$ as the current update, analogous to the gradient in CGD, and $\gamma_{\ell}\mathbf{R}_{\ell-1}$ to refine the previous search direction, corresponding to the recursive update of $\mathbf{s}_{n}$ in CGD.
The recursive update for $\mathbf{R}_{\ell}$ thus mimics $\mathbf{s}_{n}$, the search direction in CGD. The update for $\mathbf{Z}_{\ell+1}$ uses $\mathbf{R}_{\ell}$, scaled by $\alpha_{\ell}$, similarly to how CGD iterates are updated using $\mathbf{s}_{n}$. With $\mathbf{A}_{\ell}=\mathbf{I}$, this process matches CGD applied to the loss $R_{\mathbf{w}^{*}}(\mathbf{w})$ (2.11), using both current and previous gradients to refine the search direction. $\square$
3.2 Implementing $k$ Steps of LFOM with Memory Registers
We extend our analysis to show how Transformers can simulate $k$ steps of Linear First-Order Methods (LFOMs). This is achieved by maintaining a memory register at each layer, which stores accumulated updates from previous layers, simulating iterative optimization.
Proposition 2.
A memory-augmented Transformer implements $k$ steps of LFOM by maintaining memory registers across layers, where the update rules are:
$\mathbf{R}_{\ell}$ $=\mathrm{Attn}_{P_{\ell},Q_{\ell}}(\mathbf{Z}_{\ell}),$ (3.3)
$\mathbf{Z}_{\ell+1}$ $=\mathbf{Z}_{\ell}+\frac{1}{n}\sum_{j=0}^{\ell}\Gamma_{j}^{\ell}\odot\mathbf{R}_{j},$ (3.4)
and where $\Gamma_{j}^{\ell}$ governs the contribution of previous layers, and $\odot$ is the Hadamard product for scaling.
Proof Sketch.  Memformers with this architecture simulate iterative optimization by refreshing the memory register $\mathbf{R}_{\ell}$ at each layer with $\mathrm{Attn}_{P_{\ell},Q_{\ell}}(\mathbf{Z}_{\ell})$, capturing the current update. The cumulative update to $\mathbf{Z}_{\ell+1}$ incorporates past layers through a weighted sum of previous memory registers $\mathbf{R}_{j}$, with weights $\Gamma_{j}^{\ell}\in\mathbb{R}^{(d+1)\times(n+1)}$, mimicking LFOM’s cumulative iterative process.
The Hadamard product $\odot$ modulates the influence of $\mathbf{R}_{j}$, analogous to gradient preconditioning. This setup subsumes the case of diagonal preconditioners $\Lambda_{j}^{k}$ acting on gradients $\nabla R_{\mathbf{w}^{*}}(\mathbf{w}_{j}^{\mathrm{gd}})$, which in the general form looks like:
$\mathbf{w}_{k+1}^{\mathrm{gd}}=\mathbf{w}_{0}+\sum_{i=0}^{k}\Lambda_{i}^{k}\nabla R_{\mathbf{w}^{*}}(\mathbf{w}_{i}^{\mathrm{gd}}).$ (3.5)
The matrices $\Gamma_{j}^{\ell}\in\mathbb{R}^{(d+1)\times(n+1)}$ and $\Lambda_{i}^{k}\in\mathbb{R}^{d\times d}$ serve similar roles, but their dimensions differ. We expect this Hadamard product memory architecture to be able to perform richer algorithms than LFOMs, though a formal characterization of its full potential remains to be done.
The full proof follows from the cumulative memory structure and the connection between attention and preconditioned gradients, as discussed in the proof steps of Lemma 2.11. $\square$
Remark. The update in (3.4) can be interpreted as a form of "Gated Memory," analogous to gating mechanisms in LSTMs and GRUs, which also utilize the Hadamard product to regulate information flow through gates. This resemblance suggests that insights from these architectures may help refine memory mechanisms in Transformers, potentially improving their ability to handle long-term dependencies in optimization tasks. However, further investigation is needed to fully understand this connection.
If the $\Lambda_{i}^{k}$ terms in (3.5) are allowed to be general matrices and the product $\Lambda_{i}^{k}\nabla R_{\mathbf{w}^{*}}(\mathbf{w}_{i}^{\mathrm{gd}})$ is replaced by the Hadamard product $\Lambda_{i}^{k}\odot\nabla R_{\mathbf{w}^{*}}(\mathbf{w}_{i}^{\mathrm{gd}})$, we obtain a class of algorithms that sits between first-order and second-order optimization methods. In fact, this is the type of algorithm implemented by the forward pass of the Memformer under the architecture defined in (3.4). A challenging task for future explorations would be to extract a novel high-level optimization algorithm from the trained Transformer/Memformer; doing so, in a rigorous manner remains out of current reach.
3.3 Experimental Results: Memformers implementing LFOMs
In this section, we present our empirical results for Memformers implementing conjugate gradient descent (CGD), general linear first-order methods (LFOMs), and general LFOMs with $\mathrm{GD}^{++}$. The method $\mathrm{GD}^{++}$ is a quasi-Newton method where the inverse Hessian in Newton’s method is approximated by a truncated Neumann series; for more details on $\mathrm{GD}^{++}$, refer to Section A.10 of Von Oswald et al. (2023a).
We consider the in-context loss function (2.11) for linear regression. The input dimension is set to $d=5$, and the number of training observations in the prompt is $n=20$. Both the inputs $\mathbf{x}^{(i)}$ and the target weight vector $\mathbf{w}^{*}$ are sampled from Gaussian distributions: $\mathbf{x}^{(i)}\sim\mathcal{N}(0,\mathbf{\Sigma})$ and $\mathbf{w}^{*}\sim\mathcal{N}(0,\mathbf{\Sigma}^{-1})$, where $\mathbf{\Sigma}=\mathbf{U}^{\top}\mathbf{D}\mathbf{U}$. Here, $\mathbf{U}$ is a uniformly random orthogonal matrix, and $\mathbf{D}$ is a fixed diagonal matrix with entries $\operatorname{diag}(1,1,0.5,0.25,1)$.
We optimize the function $f$ (2.7) for a three-layer linear transformer using the ADAM optimizer. The matrices $\mathbf{A}_{0}$, $\mathbf{A}_{1}$, and $\mathbf{A}_{2}$ (as in (2.8)) are initialized with independent and identically distributed (i.i.d.) Gaussian entries. Each gradient step is computed using a batch of size 1000, and we resample the batch every 100 steps. We clip the gradient of each matrix to have a maximum norm of 0.01. All plots are averaged over five runs, each with a different randomly sampled $\mathbf{U}$ (and thus different $\mathbf{\Sigma}$).
Figure 1 illustrates the implementation of a CGD-like algorithm under the architecture given by (3.2). In Figure 1(a), the line-search parameters $\alpha_{\ell}$ and deflection parameters $\gamma_{\ell}$ for each layer $\ell$ are obtained by training using ADAM. By “CGD-like,” we mean that upon training the memformer using ADAM, the memformer layers learn general parameters $\alpha_{\ell}$ and $\gamma_{\ell}$ which, while they may not match the exact CGD parameters for individual observations, perform well enough on each observation to be comparable to, if not competitive with, CGD. We further explain the important issue of learning general parameters in Section 4.
Figure 1(b) presents the same experiment as Figure 1(a), using the architecture in (3.2), but with the parameters $\mathbf{A}_{\ell}$ for each layer not restricted to scalars. Thus, past gradients are accounted for, similar to CGD, but with preconditioners $\mathbf{A}_{\ell}$. This is therefore not a “CGD-like” algorithm. We aim to demonstrate that once we allow preconditioned gradients, a memformer implements a certain LFOM that distinctly outperforms CGD.
Figure 2 presents the performance of LFOM under the architecture in (3.4), where the matrix parameters $\Gamma_{j}$ for each layer $j$ are obtained by training using ADAM. In our experiments, we consider the special case of $\Gamma_{j}^{\ell}=\Gamma_{j}\ \forall\ell$, which is more natural, if we consider that each layer $j$ of the memformer has an associated $\Gamma_{j}$. Figure 2(a) shows the results on non-isotropic data, and Figure 2(b) shows the results on isotropic data. Note that this algorithm is quite similar in nature to the previous case in Figure 1(b). Here, the $\Gamma_{j}$’s essentially act as preconditioners of the gradients computed in each layer. Consequently, the graphs of Figures 1(b) and 2(a) are nearly identical. In the isotropic data experiment (Figure 2(b)), we observe that the memformer does not perform better than a linear transformer. In quadratics with isotropic data, there is no significant variation in curvature across directions; thus, incorporating past gradients via momentum offers little advantage. Momentum is more beneficial in cases with non-isotropic data.
Figure 3 presents LFOM with $\mathrm{GD}^{++}$ under the architecture in (3.4), where the $\mathbf{B}_{\ell}$ blocks in the $P_{\ell}$ matrices for each layer $\ell$ (2.8) are allowed to be non-zero. Once again, the matrix parameters for each layer $\ell$ are obtained by training using ADAM. In this case, the $\mathbf{B}_{\ell}$ matrices resemble a heavily truncated Neumann series of the inverse $\mathbf{X}\mathbf{X}^{\top}$ (2.12), resulting in a quasi-Newton method. The experiments are conducted on both non-isotropic data (Figure 3(a)) and isotropic data (Figure 3(b)).
(a) Without Preconditioning
(b) With Preconditioning
Figure 1: Comparison of linear transformer and memformer with general CGD-like parameters to actual CGD running separately on each test observation.
(a) Non-Isotropic Data
(b) Isotropic Data
Figure 2: LFOM performance on non-isotropic vs. isotropic test data (Pre = with non-trivial preconditioners). Test data is independently sampled from the same distribution as the training data.
(a) Non-Isotropic Data
(b) Isotropic Data
Figure 3: LFOM GD++ performance on non-isotropic vs. isotropic test data (Pre = with non-trivial preconditioners). Test data is independently sampled from the same distribution as the training data.
4 Experiments: Memformer LFOM Performance vs CGD
We emphasize here that the results presented in Section 3.3 compare the performance of Transformers and Memformers (which learn generic parameters) against CGD that runs on fresh observations of batch size $B=1000$, independently resampled from the same distribution. But unlike CGD that computes specific parameters for each observation, the Transformer and Memformer models learn shared parameters $P_{\ell}$, $Q_{\ell}$ (and $\alpha_{\ell}$, $\gamma_{\ell}$, or $\Gamma_{\ell}$) for each layer $\ell$, and these parameters are applied uniformly across all 1000 observations in the batch. In contrast, CGD is executed individually on each of the 1000 observations in the batch, and the average log-loss versus layers is plotted.
The advantage of Memformers in implementing LFOMs (with matrices $\Gamma_{\ell}$ restricted to scalar multiples of the identity) becomes even more pronounced when tested on training data with small batch sizes, such as $B=1$ and $B=10$. In these scenarios, the Memformers learn parameters that significantly outperform CGD running in parallel on each of the observations in those small batches. Figure 4 demonstrates this comparison.
We further provide an experimental comparison of Memformer implementing LFOM vs. Nesterov Accelerated Gradient Method and Momentum GD in the Appendix.
(a) Batch Size $B=1$
(b) Batch Size $B=10$
Figure 4: Memformer with scalar preconditioners vs. CGD performance on small batch training data ($B=1$ and $B=10$). The memformer demonstrates superior performance on the training data.
5 Experiments: Impact of Using Multi-Headed Attention
Our experiments demonstrate that increasing the number of attention heads leads to better test loss performance. Transformers with multiple attention heads improve test loss by learning diverse preconditioning matrices, allowing them to better adapt to varying data covariance structures. In our architecture, as described in (3.1), the attention values from each head are summed before being assigned to the memory register $\mathbf{R}_{\ell}$ at each layer $\ell$.
Each head captures different aspects of the data, estimating gradients from multiple perspectives. This results in faster convergence, as the heads function like an ensemble, combining different gradient estimates to improve optimization. Additionally, combining gradient estimates from multiple heads acts as implicit regularization by smoothing the gradient updates, stabilizing the learning process, and preventing overfitting. This leads to better generalization on test data.
Figure 5 compares the performance of models with 1-head and 5-head attention. The results illustrate the benefits of using multiple attention heads on convergence speed and test loss performance.
(a) 1-Head Attention Performance
(b) 5-Head Attention Performance
Figure 5: Comparison of memformer (with scalar preconditioners) performance using 1-head and 5-head attention, relative to CGD.
6 Discussion and Future Work
This work demonstrates the capability of memory-augmented Transformers (Memformers) to implement a broad range of first-order optimization methods, opening several research directions. We briefly comment on some of these aspects below.
  1. (i)
    Architectural Flexibility: Small modifications, such as (gated) memory registers, significantly enhance Transformers’ ability to learn and implement diverse optimization algorithms. Future research could explore further architectural innovations to unlock even greater capabilities.
  2. (ii)
    General Function Classes: While our approach successfully makes Transformers implement LFOMs on quadratic functions, future work should extend this to more general objective functions. Doing so may require novel training strategies, and possibly architectural adjustments to handle non-quadratic functions. The role of nonlinear attention and the MLP component of Transformers may also prove to be useful here.
  3. (iii)
    Efficiency vs. Generalization: Attention-based methods require more computation than directly implementing conjugate gradient descent or momentum GD. However, Transformers excel in learning general parameters, enabling LFOMs to generalize across new data without needing per-instance optimization. Exploring practical use of such “learned optimizers” to either warmstart a solver, or to potentially even bypass it, is a tantalizing research topic.
  4. (iv)
    Theoretical Foundations and Convergence Analysis: Strengthening the theoretical basis of Transformers’ optimization capabilities, including convergence analysis and their alignment with classical optimization theory, is another important direction for future research.
  5. (v)
    Meta-learning and Transfer Learning: The ability of Transformers to learn and generalize optimization algorithms offers exciting potential for meta-learning and transfer learning, providing new opportunities in areas where traditional optimization methods fall short.
6.1 Limitations
We briefly remark on some limitations of our current framework. For instance, while LFOMs are quite versatile, our experiments (Figure 1) indicate they do not radically outperform preconditioned GD on general quadratic problems as in (2.11), where the preconditioner matrix $\Gamma_{\ell}$ (and likewise, $\mathbf{A}_{\ell}$) for the current layer $\ell$ is the main contributor to loss performance at each update step $\ell$ (3.1). On the other hand, this behavior is likely due to the task being quadratic, and a future study that tackles more general ICL formulations will likely shed light here.
Transformers can implement second-order methods like Newton’s method (Fu et al., 2023; Giannou et al., 2024), which typically outperform LFOMs in convergence speed and accuracy. However, we reiterate that the main focus of our paper is to explore the space of first-order optimization algorithms that augmented Transformers can learn, as opposed to looking for “the best” algorithm.
7 Reproducibility Statement
We believe the following points provide a clear path for replicating our results:
  • Code Availability: The code for our experiments, including Memformers and LFOM implementations, is available at https://anonymous.4open.science/r/ICLR-2025-Memformer_LFOM.
  • Experiment Setup: Detailed descriptions of the training setup, model architecture, parameter initialization, and optimization methods are included in Sections 2 and 3.3.
  • Random Seeds: Random seeds were fixed across all experiments to ensure consistency, and they are provided in the code repository for replication.
  • Hardware Requirements: All experiments were conducted on NVIDIA T4 GPUs in Google Colab.
8 Acknowledgements
This work was conducted as part of a research project under the supervision of Prof. Suvrit Sra during Sanchayan Dutta’s visit to TU Munich. Sanchayan Dutta gratefully acknowledges the hospitality of the TU Munich School of Computation, Information and Technology. Suvrit Sra acknowledges generous support from the Alexander von Humboldt Foundation. We extend our thanks to Prof. Xiang Cheng (Duke University, USA) for his initial suggestion to explore memory-augmented approaches for implementing conjugate gradient descent (CGD) and general Linear First-Order Methods (LFOM) with linear transformers. We are also thankful to Prof. Cheng for his valuable advice and assistance with sanity checks throughout our experiments.
Supplementary Material
Appendix A Proof of Lemma 1: Equivalence to Preconditioned Gradient Descent
This proof already exists in the literature, for instance, in Subsection C.1 of Ahn et al. (2024). However, we repeat it here, to make this paper as self-contained as possible.
Consider a set of fixed samples $\mathbf{x}^{(1)},\dots,\mathbf{x}^{(n)}$, along with a fixed vector $\mathbf{w}^{*}$. Let $P=\{P_{i}\}_{i=0}^{k}$ and $Q=\{Q_{i}\}_{i=0}^{k}$ represent fixed weights, and let $\mathbf{Z}_{i}$ evolve as per equation (2.5). Define $\mathbf{X}_{i}$ as the first $d$ rows of $\mathbf{Z}_{k}$ (under equation (2.8), we have $\mathbf{X}_{i}=\mathbf{X}_{0}$ for all $i$), and let $\mathbf{Y}_{i}$ be the $(d+1)$-th row of $\mathbf{Z}_{i}$. Now, let $g(\mathbf{x},\mathbf{y},k):\mathbb{R}^{d}\times\mathbb{R}\times\mathbb{Z}\to\mathbb{R}$ be a function such that for $\mathbf{x}_{n+1}=\mathbf{x}$ and $\mathbf{y}_{n+1}^{(0)}=\mathbf{y}$, the function is defined as $g(\mathbf{x},\mathbf{y},k):=\mathbf{y}_{n+1}^{(k)}$. It’s worth noting that $\mathbf{y}_{n+1}^{(k)}=[\mathbf{Y}_{k}]_{n+1}$.
We can verify that, under equation (2.8), the update rule for $\mathbf{y}_{n+1}^{(k)}$ is given by:
$\mathbf{Y}_{k+1}=\mathbf{Y}_{k}-\frac{1}{n}\mathbf{Y}_{k}M\mathbf{X}_{0}^{\top}A_{k}\mathbf{X}_{0},$
where $M$ is a mask matrix of the form:
$M=\begin{bmatrix}I&0\\ 0&0\end{bmatrix}.$
The following points can be verified:
1. $g(\mathbf{x},\mathbf{y},k)=g(\mathbf{x},0,k)+\mathbf{y}$. To see this, note that for each $i\in\{1,\dots,n\}$, we have:
$\mathbf{y}^{(i)}_{k+1}=\mathbf{y}^{(i)}_{k}-\frac{1}{n}\sum_{j=1}^{n}\mathbf{x}^{(i)\top}A_{k}\mathbf{x}^{(j)}\mathbf{y}^{(j)}_{k}.$
Thus, $\mathbf{y}^{(i)}_{k}$ does not depend on $\mathbf{y}_{n+1}^{(t)}$ for any $t$. For $\mathbf{y}_{n+1}^{(k)}$, the update becomes:
$\mathbf{y}_{n+1}^{(k+1)}=\mathbf{y}_{n+1}^{(k)}-\frac{1}{n}\sum_{j=1}^{n}\mathbf{x}_{n+1}^{\top}A_{k}\mathbf{x}^{(j)}\mathbf{y}^{(j)}_{k},$
which clearly shows that the dependence on $\mathbf{y}_{n+1}^{(k)}$ is additive. Through a simple induction, we can establish:
$g(\mathbf{x},\mathbf{y},k+1)-\mathbf{y}=g(\mathbf{x},\mathbf{y},k)-\mathbf{y}.$
2. The function $g(\mathbf{x},0,k)$ is linear in $\mathbf{x}$. To see this, note that for $j\neq n+1$, $\mathbf{y}_{j}^{(k)}$ does not depend on $\mathbf{x}_{n+1}^{(t)}$ for any $t$, $j$, or $k$. Therefore, the update for $\mathbf{y}_{n+1}^{(k+1)}$ depends linearly on $\mathbf{x}_{n+1}$ and $\mathbf{y}_{n+1}^{(k)}$. Since $\mathbf{y}_{n+1}^{(0)}=0$ is linear in $\mathbf{x}$, we conclude by induction that the result holds.
Considering these points, we can confirm that for each $k$, there exists a vector $\theta_{k}\in\mathbb{R}^{d}$ such that:
$g(\mathbf{x},\mathbf{y},k)=g(\mathbf{x},0,k)+\mathbf{y}=\langle\theta_{k},\mathbf{x}\rangle+\mathbf{y},$
for all $\mathbf{x}$ and $\mathbf{y}$. It follows that $g(\mathbf{x},\mathbf{y},0)=\mathbf{y}$, so that $\langle\theta_{0},\mathbf{x}\rangle=g(\mathbf{x},\mathbf{y},0)-\mathbf{y}=0$, implying $\theta_{0}=0$.
We now focus on the third key fact: for each $i$, we have:
$g(\mathbf{x}^{(i)},\mathbf{y}^{(i)},k)=\mathbf{y}^{(i)}_{k}=\langle\theta_{k},\mathbf{x}^{(i)}\rangle+\mathbf{y}^{(i)}.$
To prove this, let $\mathbf{x}_{n+1}:=\mathbf{x}^{(i)}$ for some $i\in\{1,\dots,n\}$. Then:
$\mathbf{y}^{(i)}_{k+1}=\mathbf{y}^{(i)}_{k}-\frac{1}{n}\sum_{j=1}^{n}\mathbf{x}^{(i)\top}A_{k}\mathbf{x}^{(j)}\mathbf{y}^{(j)}_{k},$
$\mathbf{y}_{n+1}^{(k+1)}=\mathbf{y}_{n+1}^{(k)}-\frac{1}{n}\sum_{j=1}^{n}\mathbf{x}_{n+1}^{\top}A_{k}\mathbf{x}^{(j)}\mathbf{y}^{(j)}_{k},$
therefore, $\mathbf{y}^{(i)}_{k+1}=\mathbf{y}_{n+1}^{(k+1)}$ when $\mathbf{y}^{(i)}_{k}=\mathbf{y}_{n+1}^{(k)}$. This completes the induction, given that $\mathbf{y}^{(i)}_{0}=\mathbf{y}_{n+1}^{(0)}$ by definition.
Let $\mathbf{\bar{X}}\in\mathbb{R}^{d\times n}$ be the matrix whose columns are $\mathbf{x}^{(1)},\dots,\mathbf{x}^{(n)}$, excluding $\mathbf{x}_{n+1}$, and let $\mathbf{\bar{Y}}_{k}\in\mathbb{R}^{1\times n}$ be the vector of $\mathbf{y}^{(1)}_{k},\dots,\mathbf{y}^{(n)}_{k}$. It follows that:
$\mathbf{\bar{Y}}_{k}=\mathbf{\bar{Y}}_{0}+\theta_{k}^{\top}\mathbf{\bar{X}}.$
Using this, the update formula for $\mathbf{y}_{n+1}^{(k)}$ becomes:
$\mathbf{y}_{n+1}^{(k+1)}=\mathbf{y}_{n+1}^{(k)}-\frac{1}{n}\langle A_{k}\mathbf{\bar{X}}^{\top}\mathbf{\bar{Y}}_{k},\mathbf{x}_{n+1}\rangle,$
leading to the update:
$\langle\theta_{k+1},\mathbf{x}_{n+1}\rangle=\langle\theta_{k},\mathbf{x}_{n+1}\rangle-\frac{1}{n}\langle A_{k}\mathbf{\bar{X}}(\mathbf{\bar{X}}^{\top}\theta_{k}+\mathbf{\bar{Y}}_{0}),\mathbf{x}_{n+1}\rangle.$
Since $\mathbf{x}_{n+1}$ is arbitrary, we derive the general update formula:
$\theta_{k+1}=\theta_{k}-\frac{1}{n}A_{k}\mathbf{\bar{X}}\mathbf{\bar{X}}^{\top}(\theta_{k}+\mathbf{w}^{*}).$
Treating $A_{k}$ as a preconditioner, and letting $f(\theta):=\frac{1}{2n}(\theta+\mathbf{w}^{*})^{\top}\mathbf{\bar{X}}\mathbf{\bar{X}}^{\top}(\theta+\mathbf{w}^{*})$, we can express the update as:
$\theta_{k+1}=\theta_{k}-\frac{1}{n}A_{k}\nabla f(\theta).$
Finally, let $\mathbf{w}_{k}^{\text{gd}}:=-\theta_{k}$. We can verify that $f(-\mathbf{w})=R_{\mathbf{w}^{*}}(\mathbf{w})$, implying that:
$\mathbf{w}_{k+1}^{\text{gd}}=\mathbf{w}_{k}^{\text{gd}}-\frac{1}{n}A_{k}\nabla R_{\mathbf{w}^{*}}(\mathbf{w}_{k}^{\text{gd}}).$
We also confirm that for any $\mathbf{x}_{n+1}$, the prediction of $\mathbf{y}_{n+1}^{(k)}$ is:
$g(\mathbf{x}_{n+1},\mathbf{y}_{n+1},k)=\mathbf{y}_{n+1}-\langle\theta,\mathbf{x}_{n+1}\rangle=\mathbf{y}_{n+1}+\langle\mathbf{w}_{k}^{\text{gd}},\mathbf{x}_{n+1}\rangle.$
This concludes the proof. We have simply followed the update rule (2.5) to its logical conclusion.
Appendix B Comparison to Nesterov Accelerated Gradient Method (NAG) and Momentum Gradient Descent (MGD)
B.1 Nesterov Accelerated Gradient Method (NAG)
NAG is a commonly used optimization technique that builds on classical gradient descent by incorporating a momentum term that anticipates the next update. Specifically, the weights are updated using the following update rules:
$\mathbf{v}_{k+1}=\mathbf{w}_{k}+\beta_{k}(\mathbf{w}_{k}-\mathbf{w}_{k-1})$
$\mathbf{w}_{k+1}=\mathbf{v}_{k+1}-\eta_{k}\nabla f(\mathbf{v}_{k+1})$
Here, $\beta_{k}$ controls the influence of previous updates (momentum), and $\eta_{k}$ is the learning rate. In our experiments, we selected $\eta_{k}=0.03$ and $\beta_{k}=0.9$ after testing various values of these parameters on the given distribution, as in Section 3.3. These values provided the best performance. The momentum term allows NAG to “look ahead" in the optimization trajectory, which often leads to faster convergence than vanilla gradient descent.
B.2 Momentum Gradient Descent (MGD)
Momentum Gradient Descent operates similarly to NAG but without the anticipation of future steps. The algorithm updates the weights based on a momentum term that accelerates convergence in directions with consistent gradients. The update rule for MGD is given by:
$\mathbf{v}_{k+1}=\beta_{k}\mathbf{v}_{k}-\eta_{k}\nabla f(\mathbf{w}_{k})$
$\mathbf{w}_{k+1}=\mathbf{w}_{k}+\mathbf{v}_{k+1}$
In our experiments, the learning rate $\eta_{k}=0.005$ and momentum parameter $\beta_{k}=0.9$ provided the best results on the given distribution, as in Section 3.3. Momentum helps to mitigate oscillations in directions with high curvature, stabilizing the optimization trajectory and leading to faster convergence compared to gradient descent.
B.3 Memformer LFOMs vs. NAG and MGD
In our experiments, we observed that memformer LFOMs outperform both NAG and MGD on non-isotropic data. Figures 6(a) and 6(b) compare the performance of memformer LFOMs with NAG and MGD, respectively, on the same non-isotropic data. As shown, the memformer achieves faster convergence and much better loss performance compared to both algorithms.
(a) Nesterov AGM vs. Memformer LFOM on non-isotropic data.
(b) Momentum GD vs. Memformer LFOM on non-isotropic data.
Figure 6: Comparison of Nesterov Accelerated Gradient Method (left) and Momentum Gradient Descent (right) vs. Memformer LFOM on non-isotropic data.

### v3

@@
-Memory-augmented Transformers can implement Linear First-Order Optimization Methods
+Toward generalizable learning of all (linear) first-order methods via memory augmented Transformers
 Abstract
-We show that memory-augmented Transformers (Memformers) can implement linear first-order optimization methods such as conjugate gradient descent, momentum methods, and more generally, methods that linearly combine past gradients.
-Building on prior work that demonstrates how Transformers can simulate preconditioned gradient descent, we provide theoretical and empirical evidence that Memformers can learn more advanced optimization algorithms.
-Specifically, we analyze how memory registers in Memformers store suitable intermediate attention values allowing them to implement algorithms such as conjugate gradient.
-Our results show that Memformers can efficiently learn these methods by training on random linear regression tasks, even learning methods that outperform conjugate gradient.
-This work extends our knowledge about the algorithmic capabilities of Transformers, showing how they can learn complex optimization methods.
+We show that memory-augmented Transformers can implement the entire class of linear first-order methods (LFOMs), a class that contains gradient descent (GD) and more advanced methods such as conjugate gradient descent (CGD), momentum methods and all other variants that linearly combine past gradients.
+Building on prior work that studies how Transformers simulate GD, we provide theoretical and empirical evidence that memory-augmented Transformers can learn more advanced algorithms.
+We then take a first step toward turning the learned algorithms into actually usable methods by developing a mixture-of-experts (MoE) approach for test-time adaptation to out-of-distribution (OOD) samples.
+Lastly, we show that LFOMs can themselves be treated as learnable algorithms, whose parameters can be learned from data to attain strong performance.
 1 Introduction
-In-context learning (ICL) allows large language models (LLMs) to generate contextually appropriate outputs based solely on examples and queries provided in a prompt, without requiring any parameter adjustments (Brown, 2020; Liu et al., 2021; Lu et al., 2021; Wei et al., 2022; Wu et al., 2022).
-This remarkable ability has spurred research into understanding how Transformers can implement algorithms (Brown, 2020; Achiam et al., 2023; Touvron et al., 2023), with recent studies focusing on their capability to simulate optimization algorithms (Dai et al., 2022; Von Oswald et al., 2023a; Garg et al., 2022; Akyürek et al., 2022).
+In-context learning (ICL) allows large language models (LLMs) to generate contextually appropriate outputs based solely on examples and queries provided in a prompt (Brown, 2020; Liu et al., 2021; Lu et al., 2021; Wei et al., 2022; Wu et al., 2022).
+This remarkable ability has spurred research into understanding how Transformers can implement algorithms (Achiam et al., 2023; Touvron et al., 2023), with recent studies focusing on they can simulate optimization algorithms (Dai et al., 2022; Von Oswald et al., 2023a; Garg et al., 2022; Akyürek et al., 2022).
 Transformers have been shown to implement gradient-based optimization during their forward pass, such as preconditioned gradient descent for linear regression tasks (Dai et al., 2022; Mahankali et al., 2023; Ahn et al., 2024).
-More recently, studies have demonstrated that Transformers can learn even more advanced optimization methods.
-For instance, Fu et al.
-(2023) showed that Transformers exhibit convergence rates comparable to Iterative Newton’s Method, a higher-order optimization technique that converges exponentially faster than gradient descent for in-context linear regression.
-Additionally, Vladymyrov et al.
-(2024) proved that Transformers can, in fact, learn a variant of gradient descent that approximates second-order methods, such as $\mathrm{GD}^{++}$, achieving convergence rates similar to Newton’s method.
+Other recent work shows that Transformers can also learn more advanced optimization methods, e.g., Fu et al.
+(2023) show that Transformers exhibit convergence rates comparable to Newton’s Method for ICL on linear regression.
+Vladymyrov et al.
+(2024) prove that Transformers can learn a variant of gradient descent that approximates second-order methods, such as $\mathrm{GD}^{++}$, achieving convergence rates similar to Newton’s method.
 These findings lead to the central question of our paper:
 Can Transformers efficiently learn more advanced gradient-based optimization methods?
-We hope that by addressing this question, we can reveal some of the representational power of Transformers as “algorithm learners,” and thus further motivate the use of machine learning for discovering new optimization algorithms.
-To make our question more precise, we focus on learning the entire class of gradient-based algorithms obtained by linearly combining past gradients, known as Linear First-Order Methods (LFOMs) (Goh, 2017), where the $k+1^{\text{st}}$ iterate is
-$w^{k+1}=w^{0}+\sum_{i=0}^{k}\Gamma_{i}^{k}\nabla f(w^{i}),$ (1.1)
+We address this question by investigating the representational power of Transformers as algorithm learners, and with it, contributing to the topic of machine driven algorithm discovery.
+More specifically, we focus on learning all gradient-based algorithms obtainable by linearly combining past gradients, known as Linear First-Order Methods (LFOMs) (Goh, 2017), wherein $(k+1)$-st iterate is
+$w^{k+1}=w^{0}+\sum\nolimits_{i=0}^{k}\Gamma_{i}^{k}\nabla f(w^{i}),$ (1.1)
 and where $\{\Gamma_{i}^{k}\}_{i=0}^{k}$ are diagonal matrices.
-Model (1.1) is quite general, as it includes, as special cases, standard methods such as gradient descent (GD), momentum GD, Nesterov’s accelerated gradient, conjugate gradient, and in a stochastic setting, AdaGrad, ADAM, among others.
-Our key insight for efficiently learning LFOMs is to leverage memory-augmented Transformers, known as Memformers (Wu et al., 2020; Xu et al., 2021), which retain intermediate attention values across layers.
-This memory can enable them to store past gradients, facilitating the execution of advanced first-order methods such as conjugate gradient descent and momentum methods.
-The same mechanism allows Memformers to implement more general LFOMs.
-While unconditional learning of gradient methods remains out of reach, we build on related work demonstrating that Transformers can learn gradient descent in the context of linear regression tasks (Garg et al., 2022; Akyürek et al., 2022; Von Oswald et al., 2023a; Ahn et al., 2024; Zhang et al., 2024).
-Inspired by these findings, and extending the work of (Ahn et al., 2024), we conduct a theoretical analysis of the loss landscape for memory-augmented linear Transformers that omit softmax activation (Schlag et al., 2021; Von Oswald et al., 2023a; Ahn et al., 2024).
-In summary, our main contributions are as follows:
-Main Contributions
-  1. (A)
-    Theoretical and Experimental Evidence.
-We provide theoretical justification and empirical results showing that Memformers trained on linear regression tasks learn LFOMs (and even quasi-Newton methods like $\mathrm{GD}^{++}$ (Von Oswald et al., 2023a) for novel in-context data.
-  2. (B)
-    Performance Beyond Conjugate Gradient Descent.
-We demonstrate that LFOMs learned by Memformers significantly outperform conjugate gradient descent on training data, while remaining competitive on test data, indicating good generalization performance.
-  3. (C)
-    Multi-Headed Attention Improves Performance.
-We show empirically that multi-headed attention improves Memformers’ test performance, and offer theoretical justification for why increasing attention heads enhances loss performance on test data.
-The main objective of this paper is to investigate the potential of augmented Transformers to learn optimization algorithms.
-We are not advocating for Transformers as substitutes for established optimization methods, but rather, we aim to kindle further exploration into their algorithmic capabilities.
-We believe our results contribute to understanding how augmented Transformers could help with optimization, which may ultimately lead to new and practical gradient-based algorithms.
+Iteration (1.1) is quite general: it includes as special cases standard methods such as gradient descent (GD), momentum GD, Nesterov’s accelerated gradient, conjugate gradient, and in a stochastic version, AdaGrad, ADAM, among others.
+Toward showing how Transformers can (efficiently) capture all LFOMs, our key insight is to consider memory-augmented Transformers, known as Memformers (Wu et al., 2020; Xu et al., 2021), which retain intermediate attention values across layers.
+This memory enables Memformers to store past gradients, and facilitates efficient mimicking of first-order methods such as conjugate gradient descent and momentum methods, that also use past gradients.
+One may ask: What does it mean to “learn” an optimization algorithm? For us, it refers to two key aspects:
+1.
+Expressivity.
+The Memformer can perform iterations of LFOMs in its forward pass, under a suitable choice of parameters.
+Thus, the architecture and parameterization are sufficiently expressive to simulate LFOMs.
+2.
+Trainability.
+The Memformer’s parameters can be trained on random linear regression tasks.
+Then, using these learned parameters (which are shared across all in-context data samples), the Memformer can execute “CGD-like” and “LFOM-like” iterations during a forward pass.
+(cf.
+Section 3)
+While far from unconditional learning of GD, the above two aspects of learning follow a large body of work on how Transformers learn GD in context (Garg et al., 2022; Akyürek et al., 2022; Von Oswald et al., 2023a; Ahn et al., 2024; Zhang et al., 2024).
+Inspired by these works, and extending the work of Ahn et al.
+(2024), we study “learning” by analyzing the loss landscape for memory-augmented Linear Transformers that omit softmax activations (Schlag et al., 2021; Von Oswald et al., 2023a; Ahn et al., 2024).
+1.1 Main Contributions
+  1. 1.
+    Theoretical justification that Memformers can implement LFOM iterations, including CGD.
+We provide a rigorous theoretical framework showing that Memformers, when trained on linear regression tasks, can be configured to perform iterations of any desired LFOM in their forward pass.
+  2. 2.
+    Empirical evidence that Memformers learn optimization algorithms.
+Through extensive experiments, we demonstrate that a Memformer can learn LFOMs by training on random linear regression tasks.
+Thereafter, using its learned parameters, it can solve new regression problems using its forward pass.
+Notably, on several inputs such a Memformer performs competitively against (sometimes even outpeforming) CGD, Nesterov AGM, and momentum GD.
+    This finding is potentially significant because while CGD (and several LFOMs) adapts its step-size and conjugacy parameters per sample—which translates into $2B$ parameters $\{\alpha,\gamma\}$ per iteration, for a batch of size $B$—a Memformer reuses one set of learned weights across the entire batch (see also Remark 3.3).
+  3. 3.
+    Enhanced performance through multi-headed attention and mixture-of-experts (MoE).
+We demonstrate that multi-headed attention improves both in-distribution and out-of-distribution (OOD) performance, with an MoE approach that enables each attention head to specialize in handling different data distributions.
+Additionally, we explore the performance of LFOMs when their parameters are learned directly from data, thus framing LFOMs as statistically learnable algorithms, and showcase their ability to generalize effectively across diverse in-context data samples.
+  4. 4.
+    Connections to test-time adaptation and relevance for LLMs.
+Beyond in-context adaptation, we investigate ways to update model parameters at test time to address shifting distributions, a topic with growing importance in large language models (LLMs).
+Recent work suggests that appropriately scaling test-time compute, rather than just model size, can yield stronger performance improvements (Snell et al., 2024), and we link these insights to our MoE-based Memformer approach (Section 4.2) to demonstrate how gating and memory preconditioners can adapt at inference time to handle various distribution shifts (Sun et al., 2020).
+Our objective.
+We would like to emphasize that our main objective is to investigate the potential of Memformers to learn advanced optimization algorithms in a general sense, as well as their potential to adapt to OOD optimization problems.
+We are not advocating for Transformers as replacement for established optimization methods.
+Nevertheless, we hope that in the future, machine-discovered optimization algorithms surpass hand-designed methods such as GD or CGD, at least in an average-case / distributional sense.
 1.2 Related Work
-Research on Transformers is extremely active, and we cannot hope to fully capture the breadth of the related literature. Below, we summarize the most immediately relevant topics.
 In-Context Learning.
@@
 Subsequent works have explored how these models adapt to new tasks without requiring parameter updates (Xie et al., 2021; Von Oswald et al., 2023b; Hahn and Goyal, 2023; Liu et al., 2021; Lu et al., 2021; Wei et al., 2022; Wu et al., 2022).
-This foundational research has paved the way for studies investigating how Transformers can implement specific algorithms, such as gradient-based methods.
 Gradient-Based Methods in Transformers.
 Garg et al.
-(2022) analyze the learning of gradient descent within Transformers, particularly in the context of ICL for linear functions.
-Empirical studies (Garg et al., 2022; Akyürek et al., 2022; Von Oswald et al., 2023a) have shown that Transformers can learn gradient descent after being trained on random linear regression tasks.
+(2022) analyze the learning of GD within Transformers, particularly in the context of ICL for linear functions.
+Empirical studies (Garg et al., 2022; Akyürek et al., 2022; Von Oswald et al., 2023a) have shown that Transformers can learn GD after being trained on random linear regression tasks.
 Expanding on these results, Von Oswald et al.
 (2023a); Ahn et al.
-(2024) demonstrate that Transformers can implement preconditioned gradient descent for solving linear regression problems presented in input prompts.
-Notably, these works—as well as ours—utilize Linear Transformers as discussed in (Schlag et al., 2021; Von Oswald et al., 2023a; Ahn et al., 2023).
+(2024) demonstrate that Transformers can implement preconditioned GD for solving linear regression problems presented in input prompts.
+Notably, these works, as well as ours, utilize Linear Transformers as discussed in (Schlag et al., 2021; Von Oswald et al., 2023a; Ahn et al., 2023).
 Higher-Order Optimization Methods in Transformers.
@@
 Memory-Augmented Transformers (Memformers).
-Memformers were introduced by Wu et al.
-(2020); Xu et al.
-(2021).
-These models retain intermediate attention values across layers through memory registers, enabling more complex computations and optimization methods to be learned.
-While significant progress has been made in understanding how Transformers can learn gradient descent, their potential for implementing more sophisticated LFOMs remains largely unexplored.
-Our work addresses this gap by showing how Memformers can efficiently implement a wide range of advanced first-order and quasi-second-order optimization techniques, including CGD and momentum methods, thereby pushing the boundaries of Transformer-based architectures.
+Memformers were introduced in (Wu et al., 2020; Xu et al., 2021).
+While significant progress has been made in understanding how Transformers can learn GD, their potential for learning more sophisticated methods remains largely unexplored.
+Our work addresses this gap and shows how Memformers can efficiently implement a wide range of advanced first-order and quasi-second-order methods.
 2 Background and Problem Setup
+We follow the setup of training Transformers on random instances of linear regression, following the prior works (Garg et al., 2022; Akyürek et al., 2022; Von Oswald et al., 2023a; Ahn et al., 2024).
+We largely follow the notation and formal setup of (Ahn et al., 2024), which we recall below.
 2.1 Linear Transformers on Random Linear Regression
-We follow the setup of training Transformers on random instances of linear regression, following the prior works (Garg et al., 2022; Akyürek et al., 2022; Von Oswald et al., 2023a; Ahn et al., 2024).
-We largely use the notation and formal setup of (Ahn et al., 2024), which we now proceed to recall.
 Data Distribution.
@@
 The responses are $\mathbf{y}=[\langle\mathbf{x}(1),\mathbf{w}^{*}\rangle,\dots,\langle\mathbf{x}(n),\mathbf{w}^{*}\rangle]\in\mathbb{R}^{n}$.
-Define the input matrix $\mathbf{Z}_{0}$ as:
-$\mathbf{Z}_{0}=\begin{bmatrix}\mathbf{x}(1)&\mathbf{x}(2)&\cdots&\mathbf{x}(n)&\mathbf{x}(n+1)\\ \mathbf{y}(1)&\mathbf{y}(2)&\cdots&\mathbf{y}(n)&0\end{bmatrix}\in\mathbb{R}^{(d+1)\times(n+1)},$ (2.1)
+Define the input matrix $\mathbf{Z}_{0}\in\mathbb{R}^{(d+1)\times(n+1)}$ as:
+$\mathbf{Z}_{0}=\begin{bmatrix}\mathbf{x}(1)&\mathbf{x}(2)&\cdots&\mathbf{x}(n)&\mathbf{x}(n+1)\\ \mathbf{y}(1)&\mathbf{y}(2)&\cdots&\mathbf{y}(n)&0\end{bmatrix},$ (2.1)
 where the zero corresponds to the unknown response for $\mathbf{x}(n+1)$.
@@
 where $P=W_{v}$ and $Q=W_{k}^{\top}W_{q}$. This simplified form, as shown in Ahn et al. (2024), can implement preconditioned gradient descent, and it is the one we also use.
-Transformer Architecture.
-As in the related work, we also simplify the Transformer to consider only attention layers, using $L$ layers of linear self-attention with a residual connection.
+Architecture.
+Following the related work, we also simplify the Transformer to consider only attention layers, using $L$ layers of linear self-attention with a residual connection.
 Therefore, for each layer $\ell$, the output is updated as
 $\mathbf{Z}_{\ell+1}=\mathbf{Z}_{\ell}+\frac{1}{n}\text{Attn}_{P_{\ell},Q_{\ell}}(\mathbf{Z}_{\ell}),\quad\ell=0,1,\dots,L-1.$ (2.5)
-Using updates (2.5), with the input $\mathbf{Z}_{0}$, the final transformer output is
+Using (2.5), with the input $\mathbf{Z}_{0}$, the final Transformer output is
 $\text{TF}_{L}(\mathbf{Z}_{0};\{P_{\ell},Q_{\ell}\}_{\ell=0}^{L-1})=-[\mathbf{Z}_{L}]_{(d+1),(n+1)}.$ (2.6)
 The set of parameters $\{P_{\ell},Q_{\ell}\}_{\ell=0}^{L-1}$ is then learned by minimizing the following training objective:
-$f\left(\{P_{\ell},Q_{\ell}\}_{\ell=0}^{L-1}\right)=\mathbb{E}_{(\mathbf{Z}_{0},\mathbf{w}^{*})}\left[\left(\text{TF}_{L}(\mathbf{Z}_{0})+(\mathbf{w}^{*})^{\top}\mathbf{x}(n+1)\right)^{2}\right].$ (2.7)
-Here, the scaling factor $\frac{1}{n}$ is used only for ease of notation and does not influence the expressive power of the Transformer.
+$\mathbb{E}_{(\mathbf{Z}_{0},\mathbf{w}^{*})}\left[\left(\text{TF}_{L}(\mathbf{Z}_{0})+(\mathbf{w}^{*})^{\top}\mathbf{x}(n+1)\right)^{2}\right].$ (2.7)
 We will utilize the following lemma from Ahn et al.
 (2024), which demonstrates that multi-layer Transformers simulate preconditioned gradient descent under suitable parameterization.
-We have provided the full proof of this Lemma 2.11 in the Appendix for completeness.
-$P_{\ell}=\begin{bmatrix}\mathbf{B}_{\ell}=0_{d\times d}&0\\ 0&1\end{bmatrix},\quad Q_{\ell}=-\begin{bmatrix}\mathbf{A}_{\ell}&0\\ 0&0\end{bmatrix},\quad\mathbf{A}_{\ell},\mathbf{B}_{\ell}\in\mathbb{R}^{d\times d}.$ (2.8)
-Lemma 1 (Lemma 1, Ahn et al. (2024)).
-Consider an $L$-layer linear transformer parameterized by $\mathbf{A}_{0},\dots,\mathbf{A}_{L-1}$, as in (2.8).
-Let $y_{\ell}^{(n+1)}$ be the $(d+1,n+1)$-th entry of the $\ell$-th layer output, i.e., $y_{\ell}^{(n+1)}=[\mathbf{Z}_{\ell}]_{(d+1),(n+1)}$ for $\ell=1,\dots,L$.
+We have provided the full proof of this lemma in Appendix A for completeness.
+$P_{\ell}=\begin{bmatrix}\mathbf{B}_{\ell}=0_{d\times d}&0\\ 0&1\end{bmatrix},\quad Q_{\ell}=-\begin{bmatrix}\mathbf{A}_{\ell}&0\\ 0&0\end{bmatrix}$ (2.8)
+Lemma 2.1 (Lemma 1, Ahn et al. (2024)).
+Let an $L$-layer linear transformer be parameterized by $\mathbf{A}_{0},\dots,\mathbf{A}_{L-1}$, as in (2.8). Let $y_{\ell}^{(n+1)}=[\mathbf{Z}_{\ell}]_{(d+1),(n+1)}$ for $\ell=1,\dots,L$; then,
 $y_{\ell}^{(n+1)}=-\langle\mathbf{x}^{(n+1)},\mathbf{w}_{\ell}^{\mathrm{gd}}\rangle,$ (2.9)
@@
 with the empirical least-squares loss (with $\mathbf{X}:=[\mathbf{x}^{(1)},\ldots,\mathbf{x}^{(n)}]\in\mathbb{R}^{d\times n}$):
-$R_{\mathbf{w}^{*}}(\mathbf{w}):=\frac{1}{2n}\|\mathbf{X}^{\top}\mathbf{w}-\mathbf{X}^{\top}\mathbf{w}^{*}\|^{2}=\frac{1}{2n}(\mathbf{w}-\mathbf{w}^{*})^{\top}\mathbf{X}\mathbf{X}^{\top}(\mathbf{w}-\mathbf{w}^{*}).$ (2.11)
-Note: The Hessian of $R_{\mathbf{w}^{*}}(\mathbf{w})$ is given by:
-$\nabla^{2}R_{\mathbf{w}^{*}}(\mathbf{w})=\frac{1}{n}\mathbf{X}\mathbf{X}^{\top}.$ (2.12)
-2.2 Linear First-Order Methods
-Linear First-Order Methods (LFOMs) (Goh, 2017) are a class of optimization algorithms that lineary combine past gradients for minimizing smooth objective functions.
-They iteratively update a parameter vector $\mathbf{w}$ using the gradient of the objective function.
-The general update rule is
+$R_{\mathbf{w}^{*}}(\mathbf{w}):=\frac{1}{2n}\|\mathbf{X}^{\top}\mathbf{w}-\mathbf{X}^{\top}\mathbf{w}^{*}\|^{2}.$ (2.11)
+Note.
+As observed in §C.1 of Ahn et al.
+(2024), the term $\mathrm{Attn}_{P_{\ell},Q_{\ell}}(\mathbf{Z}_{\ell})$ in (2.5) corresponds to the preconditioned gradient $\mathbf{A}_{\ell}\nabla R_{\mathbf{w}^{*}}(\mathbf{w}_{\ell}^{\mathrm{gd}})$ of (2.11) in the update for $\mathbf{w}_{\ell+1}^{\mathrm{gd}}$.
+2.2 Linear First-Order Methods (LFOMs)
+Linear First-Order Methods (LFOMs) (Goh, 2017) optimize smooth functions by iteratively updating a parameter vector $\mathbf{w}$ with current and past gradients. The general update is
 $\mathbf{w}^{k+1}=\mathbf{w}^{k}+\alpha_{k}\mathbf{d}^{k},$ (2.12)
-where $\alpha_{k}$ is the step size and $\mathbf{d}^{k}$ is the update direction, typically related to the gradient $\nabla f(\mathbf{w}^{k})$.
-Algorithms within this family differ in how they compute $\mathbf{d}^{k}$ and choose $\alpha_{k}$.
-LFOMs can be expressed in a cumulative form. For gradient descent, unrolling (2.13) we get
-$\mathbf{w}^{k+1}=\mathbf{w}^{0}-\alpha\sum\nolimits_{i=0}^{k}\nabla f(\mathbf{w}^{i}),$ (2.14)
-while common momentum methods need an additional term incorporating past gradients, yielding
-$\mathbf{w}^{k+1}=\mathbf{w}^{0}+\sum\nolimits_{i=0}^{k}\gamma_{i}^{k}\nabla f(\mathbf{w}^{i}),$ (2.15)
-where the coefficients $\gamma_{i}^{k}$ weight previous gradients.
-More advanced methods, or general LFOMs, use diagonal matrices $\Gamma_{i}^{k}$ to coordinate-wise scale each gradient component, i.e.,
-$\mathbf{w}^{k+1}=\mathbf{w}^{0}+\sum\nolimits_{i=0}^{k}\Gamma_{i}^{k}\nabla f(\mathbf{w}^{i}).$ (2.16)
-Momentum Methods and Conjugate Gradient Descent (CGD)
-Momentum methods accelerate convergence by incorporating a momentum term, modifying the gradient to account for past updates and achieving faster convergence in relevant directions.
-Conjugate Gradient Descent (CGD), on the other hand, is a first-order method optimized for quadratic minimization, serving as a benchmark for large-scale, sparse linear systems.
-After an initial steepest descent, CGD generates directions conjugate to previous ones, leading to faster convergence than standard gradient descent.
-Both are core methods within the LFOM class, summarized below:
-Momentum Methods
-1:Initialize $\mathbf{w}_{0}$, $\mathbf{v}_{0}=0$
-2:for $n=1,2,\dots$ do
-3:     Compute the gradient:
-$\nabla f(\mathbf{w}_{n})$
-4:     Update the velocity:
-$\mathbf{v}_{n}=\beta\mathbf{v}_{n-1}-\eta\nabla f(\mathbf{w}_{n})$
-5:     Update the iterate:
-$\mathbf{w}_{n+1}=\mathbf{w}_{n}+\mathbf{v}_{n}$
-6:end for
-7:$\beta$: Momentum coefficient (controls the influence of past gradients)
-8:$\eta$: Learning rate (scales the gradient step size)
-Conjugate Gradient Descent (CGD)
-1:Initialize $\mathbf{w}_{0}$, $\mathbf{s}_{0}=-\nabla f(\mathbf{w}_{0})$
-2:for $n=1,2,\dots$ do
-3:     Compute the steepest descent direction:
-$\Delta\mathbf{w}_{n}=-\nabla f(\mathbf{w}_{n})$
-4:     Compute the conjugacy coefficient:
-$\gamma_{n}=\frac{\|\nabla f(\mathbf{w}_{n})\|^{2}}{\|\nabla f(\mathbf{w}_{n-1})\|^{2}}$
-5:     Update the search direction:
-$\mathbf{s}_{n}=\Delta\mathbf{w}_{n}+\gamma_{n}\mathbf{s}_{n-1}$
-6:     Perform a line search:
-$\alpha_{n}=\arg\min_{\alpha}f(\mathbf{w}_{n}+\alpha\mathbf{s}_{n})$
-7:     Update the iterate:
-$\mathbf{w}_{n+1}=\mathbf{w}_{n}+\alpha_{n}\mathbf{s}_{n}$
-8:end for
-Momentum methods provide fast convergence by accumulating gradient history and are widely used in modern optimization.
-CGD converges in at most $N$ iterations for quadratic functions, where $N$ is the number of variables, and is effective for ill-conditioned problems.
-3 Memformers Implement LFOMs In-Context
-Building on Lemma 2.11, we demonstrate how memory-augmented Transformers can implement Linear First-Order Methods (LFOMs) like Conjugate Gradient Descent (CGD) in-context.
-As noted in Ahn et al.
-(2024) (Subsection C.1), the term $\text{Attn}_{P_{\ell},Q_{\ell}}(\mathbf{Z}_{\ell})$ in the update for $\mathbf{Z}_{\ell+1}$ corresponds to the preconditioned gradient $\mathbf{A}_{\ell}\nabla R_{\mathbf{w}^{*}}(\mathbf{w}_{\ell}^{\mathrm{gd}})$ of the in-context loss (2.11).
-3.1 Dynamic Memory for CGD-like Algorithms
-With attention simulating preconditioned gradient descent, we now explore how memory-augmented Transformers can simulate Conjugate Gradient Descent (CGD).
-The dynamic memory mechanism enables recursive refinement of search directions, closely mimicking CGD’s iterative process.
-Proposition 1.
-A memory-augmented Transformer can implement Conjugate Gradient Descent (CGD) through a dynamic memory mechanism that recursively refines search directions, where the update rules are:
+where $\alpha_{k}$ is a step size and $\mathbf{d}^{k}$ is a direction typically tied to $\nabla f(\mathbf{w}^{k})$.
+Different LFOMs vary in how $\mathbf{d}^{k}$ and $\alpha_{k}$ are chosen.
+For example:
+[…]
+[…]
+In more advanced LFOMs, the scalars $\gamma_{i}^{k}$ are replaced by diagonal matrices $\Gamma_{i}^{k}$, enabling coordinate-wise scaling.
+Conjugate Gradient Descent (CGD).
+Specialized for quadratic minimization, CGD generates update directions $\mathbf{s}_{n}$ that are conjugate to previous ones, leading to faster convergence than standard gradient descent.
+It iterates:
+$\Delta\mathbf{w}_{n}=-\nabla f(\mathbf{w}_{n}),\quad\gamma_{n}=\frac{\|\nabla f(\mathbf{w}_{n})\|^{2}}{\|\nabla f(\mathbf{w}_{n-1})\|^{2}},$ (2.13)
+$\mathbf{s}_{n}=\Delta\mathbf{w}_{n}+\gamma_{n}\mathbf{s}_{n-1},$
+then chooses a step size $\alpha_{n}$ (often via line search) and updates: $\mathbf{w}_{n+1}=\mathbf{w}_{n}+\alpha_{n}\,\mathbf{s}_{n}.$
+Both momentum GD and CGD are LFOMs.
+Momentum methods are widely adopted in modern optimization, while CGD converges in at most $N$ iterations for $N$-dimensional quadratics, and is effective for ill-conditioned problems.
+Figure 1: CGD-like Memformer (3.2) without preconditioning ($\mathbf{A}_{\ell}=\mathbf{I}$) vs.
+actual CGD running separately on each test sample.
+Test data is drawn from the same distribution as the training data.
+Figure 2: CGD-like Memformer (3.2) with preconditioning ($\mathbf{A}_{\ell}\neq\mathbf{I}$).
+This yields a more general LFOM-like scheme, often outperforming CGD, Nesterov AGM and momentum GD.
+Test data is independently drawn from the same distribution as training data.
+Figure 3: LFOM Memformer (3.4) vs. CGD, Nesterov AGM and momentum GD (Pre = non-trivial preconditioners).
+3 Memformers Can Implement LFOMs
+Memformers can “learn” LFOMs in the sense described in Section 1.
+Each layer $\ell$ has learnable parameters $\mathbf{A}_{\ell},\mathbf{B}_{\ell}$ (2.8) and either $\alpha_{\ell},\gamma_{\ell}$ (3.2) or $\Gamma_{\ell}$ (3.4).
+Theorems 3.1 and 3.2 show that with suitable parameters, Memformers can exactly implement CGD and LFOM iterations in their forward pass.
+With only a small number of learned parameters shared across a batch of in-context test data samples, Memformers can perform “CGD-like” (Section 3.1) or “LFOM-like” (Section 3.2) updates that rival or even surpass CGD.
+We refer to algorithms realized by (3.2) as “CGD-like” and those realized by (3.4) as “LFOM-like”.
+3.1 Single Dynamic Memory for CGD-like Algorithms
+Theorem 3.1.
+A memory-augmented Transformer can implement Conjugate Gradient Descent (CGD) in its forward pass via a single memory register, where:
 $\mathbf{R}_{\ell}$ $=\mathrm{Attn}_{P_{\ell},Q_{\ell}}(\mathbf{Z}_{\ell})+\gamma_{\ell}\mathbf{R}_{\ell-1},$ (3.1)
 $\mathbf{Z}_{\ell+1}$ $=\mathbf{Z}_{\ell}+\alpha_{\ell}\frac{1}{n}\mathbf{R}_{\ell},$ (3.2)
-where $\gamma_{\ell}$ and $\alpha_{\ell}$ control past update influence and step size
+and $\gamma_{\ell},\alpha_{\ell}$ control the influence of past updates and step size, respectively.
 Proof Sketch.
-CGD refines search directions using current gradients and previous directions.
-The Transformer simulates this by using $\text{Attn}_{P_{\ell},Q_{\ell}}(\mathbf{Z}_{\ell})$ as the current update, analogous to the gradient in CGD, and $\gamma_{\ell}\mathbf{R}_{\ell-1}$ to refine the previous search direction, corresponding to the recursive update of $\mathbf{s}_{n}$ in CGD.
-The recursive update for $\mathbf{R}_{\ell}$ thus mimics $\mathbf{s}_{n}$, the search direction in CGD.
-The update for $\mathbf{Z}_{\ell+1}$ uses $\mathbf{R}_{\ell}$, scaled by $\alpha_{\ell}$, similarly to how CGD iterates are updated using $\mathbf{s}_{n}$.
-With $\mathbf{A}_{\ell}=\mathbf{I}$, this process matches CGD applied to the loss $R_{\mathbf{w}^{*}}(\mathbf{w})$ (2.11), using both current and previous gradients to refine the search direction.
+Here, $\mathbf{R}_{\ell}$ tracks a single memory register across layers.
+CGD updates its search direction by combining the current gradient with the previous direction (2.13).
+The Transformer mimics this by combining $\mathrm{Attn}_{P_{\ell},Q_{\ell}}(\mathbf{Z}_{\ell})$ (analogous to the gradient) and $\gamma_{\ell}\mathbf{R}_{\ell-1}$, and updating the next state $\mathbf{Z}_{\ell+1}$ via (3.2).
+With $\mathbf{A}_{\ell}=\mathbf{I}$, this matches CGD applied to (2.11); see Appendix A for details.
 $\square$
-3.2 Implementing $k$ Steps of LFOM with Memory Registers
-We extend our analysis to show how Transformers can simulate $k$ steps of Linear First-Order Methods (LFOMs).
-This is achieved by maintaining a memory register at each layer, which stores accumulated updates from previous layers, simulating iterative optimization.
-Proposition 2.
-A memory-augmented Transformer implements $k$ steps of LFOM by maintaining memory registers across layers, where the update rules are:
+3.2 Implementing $k$ Steps of LFOM with Separate Memory Registers at Each Layer $\ell$
+Next, we show how Transformers can simulate $k$ steps of general LFOM updates (1.1) by maintaining a separate memory register at each layer $\ell$.
+Theorem 3.2.
+A memory-augmented Transformer can implement $k$ steps of LFOM in its forward pass by maintaining separate memory registers across layers $\ell$:
 $\mathbf{R}_{\ell}$ $=\mathrm{Attn}_{P_{\ell},Q_{\ell}}(\mathbf{Z}_{\ell}),$ (3.3)
 $\mathbf{Z}_{\ell+1}$ $=\mathbf{Z}_{\ell}+\frac{1}{n}\sum_{j=0}^{\ell}\Gamma_{j}^{\ell}\odot\mathbf{R}_{j},$ (3.4)
-and where $\Gamma_{j}^{\ell}$ governs the contribution of previous layers, and $\odot$ is the Hadamard product for scaling.
+where $\Gamma_{j}^{\ell}$ weights previous layer updates and $\odot$ is a Hadamard product for scaling.
+[^1] The update (3.4) can be interpreted as a form of gated memory.
+Analogies with LSTMs or GRUs (which also use Hadamard gating) suggest potential ways to refine memory usage in Transformers.
 Proof Sketch.
-Memformers with this architecture simulate iterative optimization by refreshing the memory register $\mathbf{R}_{\ell}$ at each layer with $\mathrm{Attn}_{P_{\ell},Q_{\ell}}(\mathbf{Z}_{\ell})$, capturing the current update.
-The cumulative update to $\mathbf{Z}_{\ell+1}$ incorporates past layers through a weighted sum of previous memory registers $\mathbf{R}_{j}$, with weights $\Gamma_{j}^{\ell}\in\mathbb{R}^{(d+1)\times(n+1)}$, mimicking LFOM’s cumulative iterative process.
+Here, $\mathbf{R}_{\ell}$ is a separate register at each layer.
+At layer $\ell$, $\mathbf{R}_{\ell}$ captures the current update $\mathrm{Attn}_{P_{\ell},Q_{\ell}}(\mathbf{Z}_{\ell})$.
+The final output $\mathbf{Z}_{\ell+1}$ is updated by summing over registers $\mathbf{R}_{j}$ scaled through $\Gamma_{j}^{\ell}\in\mathbb{R}^{(n+1)\times(d+1)}$.
+This resembles cumulative gradient steps of general LFOMs with diagonal preconditioners (1.1).
+A full proof is in Appendix A.
+$\square$
 The Hadamard product $\odot$ modulates the influence of $\mathbf{R}_{j}$, analogous to gradient preconditioning.
-This setup subsumes the case of diagonal preconditioners $\Lambda_{j}^{k}$ acting on gradients $\nabla R_{\mathbf{w}^{*}}(\mathbf{w}_{j}^{\mathrm{gd}})$, which in the general form looks like:
+This setup subsumes the case of diagonal preconditioners $\Lambda_{i}^{k}$ acting on gradients $\nabla R_{\mathbf{w}^{*}}(\mathbf{w}_{i}^{\mathrm{gd}})$, which in the general form looks like:
 $\mathbf{w}_{k+1}^{\mathrm{gd}}=\mathbf{w}_{0}+\sum_{i=0}^{k}\Lambda_{i}^{k}\nabla R_{\mathbf{w}^{*}}(\mathbf{w}_{i}^{\mathrm{gd}}).$ (3.5)
 The matrices $\Gamma_{j}^{\ell}\in\mathbb{R}^{(d+1)\times(n+1)}$ and $\Lambda_{i}^{k}\in\mathbb{R}^{d\times d}$ serve similar roles, but their dimensions differ.
-We expect this Hadamard product memory architecture to be able to perform richer algorithms than LFOMs, though a formal characterization of its full potential remains to be done.
-The full proof follows from the cumulative memory structure and the connection between attention and preconditioned gradients, as discussed in the proof steps of Lemma 2.11. $\square$
-Remark.
-The update in (3.4) can be interpreted as a form of "Gated Memory," analogous to gating mechanisms in LSTMs and GRUs, which also utilize the Hadamard product to regulate information flow through gates.
-This resemblance suggests that insights from these architectures may help refine memory mechanisms in Transformers, potentially improving their ability to handle long-term dependencies in optimization tasks.
-However, further investigation is needed to fully understand this connection.
-If the $\Lambda_{i}^{k}$ terms in (3.5) are allowed to be general matrices and the product $\Lambda_{i}^{k}\nabla R_{\mathbf{w}^{*}}(\mathbf{w}_{i}^{\mathrm{gd}})$ is replaced by the Hadamard product $\Lambda_{i}^{k}\odot\nabla R_{\mathbf{w}^{*}}(\mathbf{w}_{i}^{\mathrm{gd}})$, we obtain a class of algorithms that sits between first-order and second-order optimization methods.
-In fact, this is the type of algorithm implemented by the forward pass of the Memformer under the architecture defined in (3.4).
-A challenging task for future explorations would be to extract a novel high-level optimization algorithm from the trained Transformer/Memformer; doing so, in a rigorous manner remains out of current reach.
-3.3 Experimental Results: Memformers implementing LFOMs
-In this section, we present our empirical results for Memformers implementing conjugate gradient descent (CGD), general linear first-order methods (LFOMs), and general LFOMs with $\mathrm{GD}^{++}$.
-The method $\mathrm{GD}^{++}$ is a quasi-Newton method where the inverse Hessian in Newton’s method is approximated by a truncated Neumann series; for more details on $\mathrm{GD}^{++}$, refer to Section A.10 of Von Oswald et al.
-(2023a).
-We consider the in-context loss function (2.11) for linear regression.
-The input dimension is set to $d=5$, and the number of training observations in the prompt is $n=20$.
-Both the inputs $\mathbf{x}^{(i)}$ and the target weight vector $\mathbf{w}^{*}$ are sampled from Gaussian distributions: $\mathbf{x}^{(i)}\sim\mathcal{N}(0,\mathbf{\Sigma})$ and $\mathbf{w}^{*}\sim\mathcal{N}(0,\mathbf{\Sigma}^{-1})$, where $\mathbf{\Sigma}=\mathbf{U}^{\top}\mathbf{D}\mathbf{U}$.
-Here, $\mathbf{U}$ is a uniformly random orthogonal matrix, and $\mathbf{D}$ is a fixed diagonal matrix with entries $\operatorname{diag}(1,1,0.5,0.25,1)$.
-We optimize the function $f$ (2.7) for a three-layer linear transformer using the ADAM optimizer.
-The matrices $\mathbf{A}_{0}$, $\mathbf{A}_{1}$, and $\mathbf{A}_{2}$ (as in (2.8)) are initialized with independent and identically distributed (i.i.d.) Gaussian entries.
-Each gradient step is computed using a batch of size 1000, and we resample the batch every 100 steps.
-We clip the gradient of each matrix to have a maximum norm of 0.01.
-All plots are averaged over five runs, each with a different randomly sampled $\mathbf{U}$ (and thus different $\mathbf{\Sigma}$).
-Figure 1 illustrates the implementation of a CGD-like algorithm under the architecture given by (3.2).
-In Figure 1(a), the line-search parameters $\alpha_{\ell}$ and deflection parameters $\gamma_{\ell}$ for each layer $\ell$ are obtained by training using ADAM.
-By “CGD-like,” we mean that upon training the memformer using ADAM, the memformer layers learn general parameters $\alpha_{\ell}$ and $\gamma_{\ell}$ which, while they may not match the exact CGD parameters for individual observations, perform well enough on each observation to be comparable to, if not competitive with, CGD.
-We further explain the important issue of learning general parameters in Section 4.
-Figure 1(b) presents the same experiment as Figure 1(a), using the architecture in (3.2), but with the parameters $\mathbf{A}_{\ell}$ for each layer not restricted to scalars.
-Thus, past gradients are accounted for, similar to CGD, but with preconditioners $\mathbf{A}_{\ell}$.
-This is therefore not a “CGD-like” algorithm.
-We aim to demonstrate that once we allow preconditioned gradients, a memformer implements a certain LFOM that distinctly outperforms CGD.
-Figure 2 presents the performance of LFOM under the architecture in (3.4), where the matrix parameters $\Gamma_{j}$ for each layer $j$ are obtained by training using ADAM.
+We expect this memory architecture to be able to perform richer algorithms than LFOMs, though a formal characterization of its full potential remains to be done.
+Remark 3.3.
+Crucially, unlike classical methods such as CGD, which “adapt” by computing step sizes ($\alpha$) and conjugacy coefficients ($\gamma$) separately for each sample—even within the same batch—our Memformer uses a single set of learned parameters shared across all samples.
+For a test batch of size $B$, CGD thus effectively uses $2B$ parameters per iteration, whereas the Memformer has just the trained internal parameters $\{P_{\ell},Q_{\ell},\Gamma_{\ell}\}$ (3.4).
+This amounts to a meta-optimizer that can exploit “average-case” behavior over the data distribution, rather than re-optimizing each instance individually.
+3.3 Experimental Results
+We evaluate Memformers for learning CGD, general LFOMs, and LFOMs with $\mathrm{GD}^{++}$ (a quasi-Newton method approximating the Hessian inverse via a truncated Neumann series (Von Oswald et al., 2023a).
+Setup.
+We use the in-context loss (2.11) for linear regression with $d=5$ and $n=20$.
+Each run samples a random orthogonal matrix $\mathbf{U}\in\mathbb{R}^{d\times d}$ and uses $\mathbf{D}=\mathrm{diag}(1,1,\tfrac{1}{2},\tfrac{1}{4},1)$.
+With a variance scalar $\sigma^{2}$ (default $1$), we set $\mathbf{\Sigma}=\sigma^{2}\mathbf{U}^{\top}\mathbf{D}\mathbf{U}$.
+We draw inputs $\mathbf{x}^{(i)}$ i.i.d.
+from $\mathcal{N}(\mathbf{0},\mathbf{\Sigma})$, and targets $\mathbf{w}^{*}$ from $\mathcal{N}(\mathbf{0},\mathbf{\Sigma}^{-1})$.
+We optimize the in-context prediction function $f$ (2.7) using a 4-layer linear Transformer trained via ADAM, initializing parameters $\mathbf{A}_{0},\mathbf{A}_{1},\mathbf{A}_{2}$, $\mathbf{A}_{3}$ (2.8) i.i.d.
+Gaussian, with batch size $1000$ (resampled every 100 steps) and gradient clipping of $0.01$.
+Plots average over five independent runs with different $\mathbf{U}$.
+CGD-like Architecture.
+Figure 1 shows a Memformer parameterized as in (3.2), where the globally learned $\alpha_{\ell}$ and $\gamma_{\ell}$ emulate CGD’s line search and deflection.
+Unlike standard CGD, which refines these parameters individually for each sample, our Memformer uses a single set of learned scalars shared across all data.
+Despite this approximation, performance remains competitive.
+Figure 2 extends the approach by allowing non-scalar preconditioners $\mathbf{A}_{\ell}$ (2.8), thus making the method a more general “LFOM-like” strategy that can even outperform standard CGD.
+LFOM Memformer.
+Figure 3 shows a Memformer with the LFOM-like architecture (3.4), where each layer’s $\Gamma_{\ell}$ acts as a preconditioner.
 In our experiments, we consider the special case of $\Gamma_{j}^{\ell}=\Gamma_{j}\ \forall\ell$, which is more natural, if we consider that each layer $j$ of the Memformer has an associated $\Gamma_{j}$.
-Figure 2(a) shows the results on non-isotropic data, and Figure 2(b) shows the results on isotropic data.
-Note that this algorithm is quite similar in nature to the previous case in Figure 1(b).
-Here, the $\Gamma_{j}$’s essentially act as preconditioners of the gradients computed in each layer.
-Consequently, the graphs of Figures 1(b) and 2(a) are nearly identical.
-In the isotropic data experiment (Figure 2(b)), we observe that the memformer does not perform better than a linear transformer.
-In quadratics with isotropic data, there is no significant variation in curvature across directions; thus, incorporating past gradients via momentum offers little advantage.
-Momentum is more beneficial in cases with non-isotropic data.
-Figure 3 presents LFOM with $\mathrm{GD}^{++}$ under the architecture in (3.4), where the $\mathbf{B}_{\ell}$ blocks in the $P_{\ell}$ matrices for each layer $\ell$ (2.8) are allowed to be non-zero.
-Once again, the matrix parameters for each layer $\ell$ are obtained by training using ADAM.
-In this case, the $\mathbf{B}_{\ell}$ matrices resemble a heavily truncated Neumann series of the inverse $\mathbf{X}\mathbf{X}^{\top}$ (2.12), resulting in a quasi-Newton method.
-The experiments are conducted on both non-isotropic data (Figure 3(a)) and isotropic data (Figure 3(b)).
-(a) Without Preconditioning
-(b) With Preconditioning
-Figure 1: Comparison of linear transformer and memformer with general CGD-like parameters to actual CGD running separately on each test observation.
-(a) Non-Isotropic Data
-(b) Isotropic Data
-Figure 2: LFOM performance on non-isotropic vs. isotropic test data (Pre = with non-trivial preconditioners). Test data is independently sampled from the same distribution as the training data.
-(a) Non-Isotropic Data
-(b) Isotropic Data
-Figure 3: LFOM GD++ performance on non-isotropic vs. isotropic test data (Pre = with non-trivial preconditioners). Test data is independently sampled from the same distribution as the training data.
-4 Experiments: Memformer LFOM Performance vs CGD
-We emphasize here that the results presented in Section 3.3 compare the performance of Transformers and Memformers (which learn generic parameters) against CGD that runs on fresh observations of batch size $B=1000$, independently resampled from the same distribution.
-But unlike CGD that computes specific parameters for each observation, the Transformer and Memformer models learn shared parameters $P_{\ell}$, $Q_{\ell}$ (and $\alpha_{\ell}$, $\gamma_{\ell}$, or $\Gamma_{\ell}$) for each layer $\ell$, and these parameters are applied uniformly across all 1000 observations in the batch.
-In contrast, CGD is executed individually on each of the 1000 observations in the batch, and the average log-loss versus layers is plotted.
-The advantage of Memformers in implementing LFOMs (with matrices $\Gamma_{\ell}$ restricted to scalar multiples of the identity) becomes even more pronounced when tested on training data with small batch sizes, such as $B=1$ and $B=10$.
-In these scenarios, the Memformers learn parameters that significantly outperform CGD running in parallel on each of the observations in those small batches.
-Figure 4 demonstrates this comparison.
-We further provide an experimental comparison of Memformer implementing LFOM vs. Nesterov Accelerated Gradient Method and Momentum GD in the Appendix.
-(a) Batch Size $B=1$
-(b) Batch Size $B=10$
-Figure 4: Memformer with scalar preconditioners vs. CGD performance on small batch training data ($B=1$ and $B=10$). The memformer demonstrates superior performance on the training data.
-5 Experiments: Impact of Using Multi-Headed Attention
-Our experiments demonstrate that increasing the number of attention heads leads to better test loss performance.
-Transformers with multiple attention heads improve test loss by learning diverse preconditioning matrices, allowing them to better adapt to varying data covariance structures.
-In our architecture, as described in (3.1), the attention values from each head are summed before being assigned to the memory register $\mathbf{R}_{\ell}$ at each layer $\ell$.
-Each head captures different aspects of the data, estimating gradients from multiple perspectives.
-This results in faster convergence, as the heads function like an ensemble, combining different gradient estimates to improve optimization.
-Additionally, combining gradient estimates from multiple heads acts as implicit regularization by smoothing the gradient updates, stabilizing the learning process, and preventing overfitting.
-This leads to better generalization on test data.
-Figure 5 compares the performance of models with 1-head and 5-head attention. The results illustrate the benefits of using multiple attention heads on convergence speed and test loss performance.
-(a) 1-Head Attention Performance
-(b) 5-Head Attention Performance
-Figure 5: Comparison of memformer (with scalar preconditioners) performance using 1-head and 5-head attention, relative to CGD.
-6 Discussion and Future Work
-This work demonstrates the capability of memory-augmented Transformers (Memformers) to implement a broad range of first-order optimization methods, opening several research directions.
-We briefly comment on some of these aspects below.
-  1. (i)
-    Architectural Flexibility: Small modifications, such as (gated) memory registers, significantly enhance Transformers’ ability to learn and implement diverse optimization algorithms.
-Future research could explore further architectural innovations to unlock even greater capabilities.
-  2. (ii)
-    General Function Classes: While our approach successfully makes Transformers implement LFOMs on quadratic functions, future work should extend this to more general objective functions.
-Doing so may require novel training strategies, and possibly architectural adjustments to handle non-quadratic functions.
-The role of nonlinear attention and the MLP component of Transformers may also prove to be useful here.
-  3. (iii)
+In practice, these parameters further improve performance.
+Figure 4 extends LFOM with $\mathrm{GD}^{++}$ by allowing the $\mathbf{B}_{\ell}$ blocks (2.8) to be non-zero.
+In this case, the $\mathbf{B}_{\ell}$ matrices resemble a heavily truncated Neumann series of the inverse $\mathbf{X}\mathbf{X}^{\top}$ (Hessian of (2.11)).
+Our experiments with more than 4 layers are detailed in Appendix C.
+Figure 4: LFOM Memformer with GD++ (3.4) vs. CGD, where the $\mathbf{B}_{\ell}$ blocks (2.8) approximate the Hessian inverse (quasi-Newton).
+3.4 Influence of Batch Size on Performance
+The results in Section 3.3 compares the performance of Transformers and Memformers (which learn shared generic parameters during training) against CGD, which computes specific parameters for each observation in a batch of size $B=1000$, independently resampled from the same distribution.
+While CGD optimizes each observation individually, Transformers and Memformers use shared parameters $P_{\ell},Q_{\ell}$ (and $\alpha_{\ell},\gamma_{\ell}$, or $\Gamma_{\ell}$) for each layer $\ell$, uniformly used for all 1000 observations in the batch.
+The average log-loss versus layers are plotted for comparison in Figures 2 and 3.
+The strength of LFOM Memformers (3.4) (with matrices $\Gamma_{\ell}$ restricted to scalar multiples of the identity) becomes even more pronounced when trained on smaller batch sizes, such as $B=1$ and $B=10$.
+In these cases, Memformers learn parameters that significantly outperform CGD, even when CGD runs independently on each observation in the batch.
+Figures 5 and 6 illustrate this comparison.
+Figure 5: LFOM Memformer (3.4) with scalar preconditioners $\Gamma_{\ell}$ vs. CGD performance on small batch size ($B=1$). The Memformer demonstrates superior performance.
+Figure 6: LFOM Memformer (3.4) with scalar preconditioners $\Gamma_{\ell}$ vs. CGD performance on small batch size ($B=10$). The Memformer outperforms CGD on the training data.
+4 Multi-Headed Attention and Out-of-Distribution (OOD) Performance
+4.1 OOD Adaptation using Multi-Headed Attention
+Next, we explore how multi-head attention further improves both in-distribution and OOD performance.
+In our experiments, increasing the number of heads yields notable gains in test-loss metrics.
+By learning diverse preconditioning matrices, multi-head architectures help Memformers adapt to data with varying covariance structures.
+Concretely, in the update (3.1), we sum the attention outputs from different heads into a single memory register $\mathbf{R}_{\ell}$ at each layer, thus aggregating multiple “views” of the underlying data distribution.
+Figure 7: Performance of an LFOM Memformer (3.4) (with scalar preconditioners $\Gamma_{\ell}$) comparing one-head and five-head attention, relative to CGD.
+Multi-head attention accelerates convergence and improves test performance.
+These findings align with prior studies underscoring the critical role of multi-head attention in in-context learning.
+For example, Chen et al.
+(2024) highlight its importance in sparse linear regression, and Cui et al.
+(2024) demonstrate its theoretical and empirical benefits.
+As illustrated in Figure 7, the number of heads strongly influences both convergence speed and final test loss.
+4.2 Test-Time Adaptation and Gated Memory
+Building on the above, we next consider test-time adaptation and show how gating mechanisms allow robust performance under distribution shift.
+Mixture of Experts (MoE) for Variance Adaptation.
+To extend multi-head architectures to OOD scenarios, we employ a mixture of experts (MoE) approach Jacobs et al.
+(1991) in a 3-head, 4-layer Memformer.
+Each head specializes in Gaussian inputs with different variances $\sigma^{2}\in\{1.0,2.0,3.0\}$.
+At inference, a learnable gating mechanism scales the layerwise output $\sum_{j=0}^{\ell}\Gamma_{j}\odot\mathbf{R}_{j}$ (3.4) per head coordinate-wise.
+Concretely, each head $H_{i}$ has a scalar gating coefficient $\alpha_{i}\in\mathbb{R}$, enabling the model to adapt effectively to unseen variances (e.g., $\sigma^{2}=\{0.5,1.5,3.5\}$).
+As shown in Figure 8, this design ensures robust generalization even when the variance differs from the training set.
+Figure 8: MoE-based 3-head, 4-layer Memformer (3.4) evaluated on Gaussian inputs of varying unseen variances. A trainable gating mechanism allows effective OOD adaptation.
+Gaussian Mixture Models (GMMs).
+We further examine OOD adaptation in a 3-head, 4-layer Memformer tested on a Gaussian Mixture Model (GMM) with different means ($\{2.5,0.0,-1.0\}$), variances ($\{0.5,1.0,2.0\}$), and mixture weights ($\{0.2,0.3,0.5\}$).
+Each head specializes in one mixture component.
+As before, a gating mechanism similarly scales the layerwise output $\sum_{j=0}^{\ell}\Gamma_{j}\odot\mathbf{R}_{j}$ across heads.
+For this GMM setup, we observe that using a 3-head, 4-layer Linear Transformer (cf.
+Figure 9) gives better performance than a 3-head, 4-layer Memformer (cf.
+Figure 10), as the $\Gamma_{j}$ matrices of each head in the Memformer overfit to the individual GMM components.
+However, this issue can be overcome, with a higher test-time budget, if the $\Gamma_{j}$ matrices are also allowed to be trainable at test time (see Figure 10).
+Figure 9: A 3-head, 4-layer Linear Transformer tested on a GMM. Each head specializes in one mixture component, and a gating module adjusts the model for OOD data.
+Figure 10: A 3-head, 4-layer Memformer (3.4) tested on a GMM.
+The gray line shows results with a learned gating scheme, while the black line also includes test-time updates to the $\Gamma_{j}$ matrices.
+Theorem 4.1 (Multi-Head Memformer with Soft Gating).
+Consider a multi-head Memformer with $H$ heads, each parameterized by $\{P_{\ell}^{h},Q_{\ell}^{h},\Gamma_{\ell}^{h}\}$.
+Suppose $\mathbf{Z}_{0}$ is drawn from a mixture of $M$ Gaussian components $\{\mathcal{N}(\mu_{m},\Sigma_{m})\}_{m=1}^{M}$, where each $\mathbf{x}^{(i)}\sim\mathcal{N}(\mu_{m},\Sigma_{m})$ and $\mathbf{w}^{*}\sim\mathcal{N}(\mu_{m},\Sigma_{m}^{-1})$.
+If each component $m$ has at least one head $h_{m}$ achieving near-optimal first-order performance, then there exist scalar gating coefficients $\alpha_{m,h}$ such that for any new prompt from component $m$, the Memformer achieves near-optimal performance via soft-gating:
+$\alpha_{m,1}\,\mathrm{Head}_{1}\;+\;\cdots\;+\;\alpha_{m,H}\,\mathrm{Head}_{H}.$
+[…]
+A formal proof of Theorem 4.1 appears in Appendix A.
+As indicated above, a Transformer can handle OOD data by fine-tuning specific test-time parameters (e.g., the gating coefficients or the diagonal of $\Gamma_{i}$, as in Figure 10).
+This design in effect operates as a meta-version of in-context learning: the main architecture parameters remain fixed, while gating and memory preconditioners can adjust to accommodate distribution shifts, depending on the test-time budget.
+5 LFOMs as Learnable Algorithms
+Finally, LFOM updates can themselves be learned as a global meta-optimizer, bridging the gap between classical algorithms and in-context algorithmic learning.
+One can directly learn the parameters of a linear first-order method (LFOM) from data, treating it as a global algorithm that generalizes across random in-context samples.
+The following theorem provides a standard finite-sample guarantee for such an approach under mild regularity conditions (boundedness, sub-Gaussianity, continuity, and compactness).
+A complete version and proof appear in Appendix A.
+Theorem 5.1 (Statistical Learnability of LFOMs in the In-Context Setting).
+Fix a class of LFOMs parameterized by $\theta\in\Theta\subset\mathbb{R}^{p}$, e.g., via diagonal preconditioners.
+Each LFOM $\theta$ maps an in-context prompt $\mathbf{Z}_{0}$ to a final prediction $\hat{y}_{\theta}(\mathbf{Z}_{0})$ for $\mathbf{x}(n+1)$, incurring the loss $\ell(\theta;\mathbf{Z}_{0})=\bigl{(}\hat{y}_{\theta}(\mathbf{Z}_{0})-\langle\mathbf{x}(n+1),\mathbf{w}^{*}\rangle\bigr{)}^{2}$.
+Drawing $N$ i.i.d.
+prompts $\{\mathbf{Z}_{0}^{(i)}\}_{i=1}^{N}$ from $\mathcal{D}$, define $\widehat{\theta}=\arg\min_{\theta\in\Theta}\frac{1}{N}\sum_{i=1}^{N}\ell\bigl{(}\theta;\mathbf{Z}_{0}^{(i)}\bigr{)}$.
+Then, with high probability over the sample draw,
+$\mathbb{E}_{\mathbf{Z}_{0}\sim\mathcal{D}}\bigl{[}\ell(\widehat{\theta};\,\mathbf{Z}_{0})\bigr{]}\;\leq\;\min_{\theta\in\Theta}\mathbb{E}_{\mathbf{Z}_{0}\sim\mathcal{D}}\bigl{[}\ell(\theta;\,\mathbf{Z}_{0})\bigr{]}\;+\;\epsilon(N,\delta),$
+where $\epsilon(N,\delta)\to 0$ as $N\to\infty$. In other words, the learned LFOM parameters $\widehat{\theta}$ generalize well on new in-context samples from $\mathcal{D}$.
+Note that our Memformer update rule (3.4) encompasses an even broader family of optimization strategies than classical LFOMs.
+Figure 11 illustrates how learning LFOM parameters (e.g.
+diagonal preconditioners) directly from data (using ADAM or similar) can match or exceed the performance of Conjugate Gradient Descent (CGD), while using one shared set of parameters across multiple samples.
+Such a strategy can overfit if batches are extremely small, but tends to generalize effectively for moderate to large batch sizes.
+Figure 11: Comparison of a learned LFOM approach vs.
+standard Conjugate Gradient Descent (CGD) on a synthetic quadratic problem (2.11).
+LFOM parameters (3.5) are trained directly from data, while CGD is run separately on each instance.
+($B$ = Batch Size)
+6 Conclusion and Future Work
+We find that memory-augmented Transformers (Memformers) can learn and implement a wide range of first-order optimization algorithms, including gradient descent, conjugate gradient, and momentum methods.
+This versatility highlights their potential as general-purpose meta-optimizers.
+Below are several avenues for further exploration:
+Architectural Extensions, MoE, and TTA.
+Additional gating strategies, mixture-of-experts (MoE), and refined memory modules may boost performance on out-of-distribution (OOD) data.
+Even a small number of experts can deliver robust gains, though identifying the “right” number of experts remains open.
+More advanced test-time adaptation (TTA), such as fine-tuning the gating or expert parameters on new data, could further enhance generalization to shifting distributions.
+Beyond Quadratic Objectives.
+Most theoretical work on algorithm learning in Transformers has focused on linear regression, leaving open the understanding of more complex, potentially nonconvex objectives.
+By integrating nonlinear attention layers or MLP components, Memformers could tackle tasks like logistic regression or PDE-based inverse problems, where simple quadratic assumptions are insufficient.
+Analyzing Memformers’ in-context learning on such domains could unify meta-learning research with more realistic, large-scale optimization problems, while clarifying theoretical guarantees beyond the linear-quadratic case.
 Efficiency vs.
-Generalization: Attention-based methods require more computation than directly implementing conjugate gradient descent or momentum GD.
-However, Transformers excel in learning general parameters, enabling LFOMs to generalize across new data without needing per-instance optimization.
-Exploring practical use of such “learned optimizers” to either warmstart a solver, or to potentially even bypass it, is a tantalizing research topic.
-  4. (iv)
-    Theoretical Foundations and Convergence Analysis: Strengthening the theoretical basis of Transformers’ optimization capabilities, including convergence analysis and their alignment with classical optimization theory, is another important direction for future research.
-  5. (v)
-    Meta-learning and Transfer Learning: The ability of Transformers to learn and generalize optimization algorithms offers exciting potential for meta-learning and transfer learning, providing new opportunities in areas where traditional optimization methods fall short.
-6.1 Limitations
-We briefly remark on some limitations of our current framework.
-For instance, while LFOMs are quite versatile, our experiments (Figure 1) indicate they do not radically outperform preconditioned GD on general quadratic problems as in (2.11), where the preconditioner matrix $\Gamma_{\ell}$ (and likewise, $\mathbf{A}_{\ell}$) for the current layer $\ell$ is the main contributor to loss performance at each update step $\ell$ (3.1).
-On the other hand, this behavior is likely due to the task being quadratic, and a future study that tackles more general ICL formulations will likely shed light here.
-Transformers can implement second-order methods like Newton’s method (Fu et al., 2023; Giannou et al., 2024), which typically outperform LFOMs in convergence speed and accuracy.
-However, we reiterate that the main focus of our paper is to explore the space of first-order optimization algorithms that augmented Transformers can learn, as opposed to looking for “the best” algorithm.
-7 Reproducibility Statement
+Generalization.
+While attention-based models can be more computationally demanding than methods like momentum or CGD, they exhibit strong generalization across diverse tasks.
+Systematic studies of runtime and accuracy trade-offs could clarify the conditions under which learned optimizers outperform classical approaches.
+Meta-Learning and Transfer.
+Memformers naturally connect to meta-learning by reusing learned optimization strategies.
+Future investigations might examine more sophisticated TTA, gating, or lightweight fine-tuning approaches to handle distribution shifts.
+Understanding how best to structure MoE—such as how many experts to deploy and how to select them on the fly—promises both practical efficiency and robust transfer in a variety of domains.
+Limitations of our framework.
+While Memformers are versatile, our experiments (Figures 2, 3) show they do not drastically outperform advanced methods on quadratic tasks (2.11), likely due to the task structure.
+Future work on broader ICL formulations may shed light on this matter.
+Notably, Transformers can implement second-order methods like Newton’s (Fu et al., 2023; Giannou et al., 2024), which typically converge faster and more accurately than LFOMs.
+However, our focus in this paper is on exploring first-order optimization algorithms that augmented Transformers can learn, rather than proposing a one-size-fits-all optimizer.
+Reproducibility Statement
 We believe the following points provide a clear path for replicating our results:
-  • Code Availability: The code for our experiments, including Memformers and LFOM implementations, is available at https://anonymous.4open.science/r/ICLR-2025-Memformer_LFOM.
+  • Code Availability: The code for our experiments, including Memformers and LFOM implementations, is available at https://anonymous.4open.science/r/ICML-2025-LFOM_Memformer.
   • Experiment Setup: Detailed descriptions of the training setup, model architecture, parameter initialization, and optimization methods are included in Sections 2 and 3.3.
@@
   • Hardware Requirements: All experiments were conducted on NVIDIA T4 GPUs in Google Colab.
-8 Acknowledgements
-This work was conducted as part of a research project under the supervision of Prof.
-Suvrit Sra during Sanchayan Dutta’s visit to TU Munich.
-Sanchayan Dutta gratefully acknowledges the hospitality of the TU Munich School of Computation, Information and Technology.
-Suvrit Sra acknowledges generous support from the Alexander von Humboldt Foundation.
-We extend our thanks to Prof.
-Xiang Cheng (Duke University, USA) for his initial suggestion to explore memory-augmented approaches for implementing conjugate gradient descent (CGD) and general Linear First-Order Methods (LFOM) with linear transformers.
-We are also thankful to Prof.
-Cheng for his valuable advice and assistance with sanity checks throughout our experiments.
 Supplementary Material
-Appendix A Proof of Lemma 1: Equivalence to Preconditioned Gradient Descent
+Appendix A Proofs
+A.1 Proof of Lemma 2.11: Equivalence to Preconditioned Gradient Descent
 This proof already exists in the literature, for instance, in Subsection C.1 of Ahn et al. (2024). However, we repeat it here, to make this paper as self-contained as possible.
@@
 We can verify that, under equation (2.8), the update rule for $\mathbf{y}_{n+1}^{(k)}$ is given by:
-$\mathbf{Y}_{k+1}=\mathbf{Y}_{k}-\frac{1}{n}\mathbf{Y}_{k}M\mathbf{X}_{0}^{\top}A_{k}\mathbf{X}_{0},$
+$\mathbf{Y}_{k+1}=\mathbf{Y}_{k}-\frac{1}{n}\mathbf{Y}_{k}M\mathbf{X}_{0}^{\top}A_{k}\mathbf{X}_{0},$ (A.1)
 where $M$ is a mask matrix of the form:
@@
 Using this, the update formula for $\mathbf{y}_{n+1}^{(k)}$ becomes:
-$\mathbf{y}_{n+1}^{(k+1)}=\mathbf{y}_{n+1}^{(k)}-\frac{1}{n}\langle A_{k}\mathbf{\bar{X}}^{\top}\mathbf{\bar{Y}}_{k},\mathbf{x}_{n+1}\rangle,$
+$\mathbf{y}_{n+1}^{(k+1)}=\mathbf{y}_{n+1}^{(k)}-\frac{1}{n}\langle A_{k}\mathbf{\bar{X}}^{\top}\mathbf{\bar{Y}}_{k},\mathbf{x}_{n+1}\rangle,$ (A.2)
 leading to the update:
-$\langle\theta_{k+1},\mathbf{x}_{n+1}\rangle=\langle\theta_{k},\mathbf{x}_{n+1}\rangle-\frac{1}{n}\langle A_{k}\mathbf{\bar{X}}(\mathbf{\bar{X}}^{\top}\theta_{k}+\mathbf{\bar{Y}}_{0}),\mathbf{x}_{n+1}\rangle.$
+$\langle\theta_{k+1},\mathbf{x}_{n+1}\rangle=\langle\theta_{k},\mathbf{x}_{n+1}\rangle-\frac{1}{n}\langle A_{k}\mathbf{\bar{X}}(\mathbf{\bar{X}}^{\top}\theta_{k}+\mathbf{\bar{Y}}_{0}),\mathbf{x}_{n+1}\rangle.$ (A.3)
 Since $\mathbf{x}_{n+1}$ is arbitrary, we derive the general update formula:
-$\theta_{k+1}=\theta_{k}-\frac{1}{n}A_{k}\mathbf{\bar{X}}\mathbf{\bar{X}}^{\top}(\theta_{k}+\mathbf{w}^{*}).$
+$\theta_{k+1}=\theta_{k}-\frac{1}{n}A_{k}\mathbf{\bar{X}}\mathbf{\bar{X}}^{\top}(\theta_{k}+\mathbf{w}^{*}).$ (A.4)
 Treating $A_{k}$ as a preconditioner, and letting $f(\theta):=\frac{1}{2n}(\theta+\mathbf{w}^{*})^{\top}\mathbf{\bar{X}}\mathbf{\bar{X}}^{\top}(\theta+\mathbf{w}^{*})$, we can express the update as:
-$\theta_{k+1}=\theta_{k}-\frac{1}{n}A_{k}\nabla f(\theta).$
+$\theta_{k+1}=\theta_{k}-\frac{1}{n}A_{k}\nabla f(\theta).$ (A.5)
 Finally, let $\mathbf{w}_{k}^{\text{gd}}:=-\theta_{k}$. We can verify that $f(-\mathbf{w})=R_{\mathbf{w}^{*}}(\mathbf{w})$, implying that:
-$\mathbf{w}_{k+1}^{\text{gd}}=\mathbf{w}_{k}^{\text{gd}}-\frac{1}{n}A_{k}\nabla R_{\mathbf{w}^{*}}(\mathbf{w}_{k}^{\text{gd}}).$
+$\mathbf{w}_{k+1}^{\text{gd}}=\mathbf{w}_{k}^{\text{gd}}-\frac{1}{n}A_{k}\nabla R_{\mathbf{w}^{*}}(\mathbf{w}_{k}^{\text{gd}}).$ (A.6)
 We also confirm that for any $\mathbf{x}_{n+1}$, the prediction of $\mathbf{y}_{n+1}^{(k)}$ is:
@@
 This concludes the proof. We have simply followed the update rule (2.5) to its logical conclusion.
+A.2 Full Proof of Theorem 3.1
+Theorem.
+A memory-augmented Transformer can implement Conjugate Gradient Descent (CGD) through a dynamic memory mechanism that recursively refines search directions, where the update rules are:
+$\mathbf{R}_{\ell}$ $=\mathrm{Attn}_{P_{\ell},Q_{\ell}}(\mathbf{Z}_{\ell})+\gamma_{\ell}\mathbf{R}_{\ell-1},$ (A.7)
+$\mathbf{Z}_{\ell+1}$ $=\mathbf{Z}_{\ell}+\alpha_{\ell}\frac{1}{n}\mathbf{R}_{\ell},$ (A.8)
+where $\gamma_{\ell}$ and $\alpha_{\ell}$ control past update influence and step size.
+Proof
+Our goal is to demonstrate that, under appropriate parameter configurations, the memory-augmented Transformer updates given by equations (A.7) and (A.8) correspond precisely to the Conjugate Gradient Descent (CGD) algorithm when applied to the quadratic loss function:
+$R_{\mathbf{w}^{*}}(\mathbf{w})=\frac{1}{2n}(\mathbf{w}-\mathbf{w}^{*})^{\top}\mathbf{X}\mathbf{X}^{\top}(\mathbf{w}-\mathbf{w}^{*}).$ (A.9)
+We will establish a mapping between the Transformer’s operations and the steps of the CGD algorithm, demonstrating that the Transformer can implement CGD under certain parameter settings.
+CGD Algorithm for Quadratic Functions
+For minimizing a quadratic function, the CGD algorithm proceeds as follows:
+{mdframed}
+[roundcorner=5pt, backgroundcolor=gray!10] Algorithm: Conjugate Gradient Descent (CGD)
+  • Initialize: $\mathbf{w}_{0}$, $\mathbf{r}_{0}=-\nabla f(\mathbf{w}_{0})$, $\mathbf{s}_{0}=\mathbf{r}_{0}$
+  • For $n=1,2,\dots$:
+    • Compute the residual:
+      $\mathbf{r}_{n}=-\nabla f(\mathbf{w}_{n})$
+    • Compute the conjugacy coefficient:
+      $\gamma_{n}=\frac{\mathbf{r}_{n}^{\top}\mathbf{r}_{n}}{\mathbf{r}_{n-1}^{\top}\mathbf{r}_{n-1}}$
+    • Update the search direction:
+      $\mathbf{s}_{n}=\mathbf{r}_{n}+\gamma_{n}\mathbf{s}_{n-1}$
+    • Compute the step size:
+      $\alpha_{n}=\frac{\mathbf{r}_{n}^{\top}\mathbf{r}_{n}}{\mathbf{s}_{n}^{\top}\mathbf{H}\mathbf{s}_{n}}$
+    • Update the parameters:
+      $\mathbf{w}_{n+1}=\mathbf{w}_{n}+\alpha_{n}\mathbf{s}_{n}$
+  • EndFor
+Mapping CGD Updates to Transformer Updates
+We first recall that in the proof of Lemma 1 (A.1), the $\mathbf{w}_{k+1}^{\text{gd}}$ update rule
+$\mathbf{w}_{k+1}^{\text{gd}}=\mathbf{w}_{k}^{\text{gd}}-\frac{1}{n}A_{k}\nabla R_{\mathbf{w}^{*}}(\mathbf{w}_{k}^{\text{gd}}),$ (A.10)
+is a direct downstream consequence of the $\mathbf{Z}_{\ell+1}$ update rule (2.5)
+$\mathbf{Z}_{\ell+1}=\mathbf{Z}_{\ell}+\frac{1}{n}\mathrm{Attn}_{P_{\ell},Q_{\ell}}(\mathbf{Z}_{\ell}),\quad\ell=0,1,\dots,L-1,$
+under the parameterization given in equation (2.8).
+Thus, the $\mathrm{Attn}_{P_{\ell},Q_{\ell}}$ term in the $\mathbf{Z}_{\ell}$ update equation is, in a precise sense, paralleled by the $-\frac{1}{n}A_{k}\nabla R_{\mathbf{w}^{*}}(\mathbf{w}_{k}^{\text{gd}})$ term in the $\mathbf{w}_{k+1}^{\text{gd}}$ update equation (A.10).
+Step 1: Initialization
+  • CGD:
+    $\mathbf{w}_{0}\text{ given},\quad\mathbf{r}_{0}=-\nabla f(\mathbf{w}_{0}),\quad\mathbf{s}_{0}=\mathbf{r}_{0}.$
+  • Transformer:
+    • The initial state $\mathbf{Z}_{0}$ in (2.5) parallels $\mathbf{w}_{0}$ in (A.10).
+    • The memory register $\mathbf{R}$ is initialized to $\mathrm{Attn}_{P_{0},Q_{0}}(\mathbf{Z}_{0})$, i.e., $\mathbf{R}_{0}=\mathrm{Attn}_{P_{0},Q_{0}}(\mathbf{Z}_{0})$, corresponding to $\mathbf{s}_{0}=\mathbf{r}_{0}$.
+    • We set $\gamma_{0}=0$, consistent with CGD initialization.
+Step 2: Update Memory Register (Search Direction)
+  • Transformer Memory Update:
+    $\mathbf{R}_{\ell}=\mathrm{Attn}_{P_{\ell},Q_{\ell}}(\mathbf{Z}_{\ell})+\gamma_{\ell}\mathbf{R}_{\ell-1}.$
+  • Correspondence with CGD:
+    $\mathbf{s}_{n}=\mathbf{r}_{n}+\gamma_{n}\mathbf{s}_{n-1}.$
+    Identifying $\mathbf{R}_{\ell}\leftrightarrow\mathbf{s}_{n}$, $\gamma_{\ell}=\gamma_{n}$, and $\mathbf{R}_{\ell-1}\leftrightarrow\mathbf{s}_{n-1}$, the Transformer’s memory update matches CGD.
+Step 3: Update Parameters
+  • Transformer Parameter Update:
+    $\mathbf{Z}_{\ell+1}=\mathbf{Z}_{\ell}+\alpha_{\ell}\frac{1}{n}\mathbf{R}_{\ell}.$
+  • Correspondence with CGD:
+    $\mathbf{w}_{n+1}=\mathbf{w}_{n}+\alpha_{n}\mathbf{s}_{n}.$
+    The scaling factor $\frac{1}{n}$ accounts for the gradient’s scaling, consistent with the CGD update when considering the Hessian $\mathbf{H}=\frac{1}{n}\mathbf{X}\mathbf{X}^{\top}$.
+Step 4: Conjugacy Coefficient $\gamma_{\ell}$ and Step Size $\alpha_{\ell}$
+  • CGD Computations: Scalar values computed based on residuals and the Hessian.
+  • Transformer Implementation:
+    • $\gamma_{\ell}$ and $\alpha_{\ell}$ are treated as parameters, ensuring structural correspondence.
+    • The Transformer’s architecture allows these as fixed or learnable parameters.
+Therefore, under suitable parameter configurations, the memory-augmented Transformer can implement CGD, demonstrating the feasibility of using the Transformer’s architecture to perform CGD-like updates.
+A.3 Full Proof of Theorem 3.2
+Theorem.
+A memory-augmented Transformer can implement $k$ steps of Linear First-Order Methods (LFOMs) by maintaining memory registers across layers, where the update rules are:
+$\mathbf{R}_{\ell}=\mathrm{Attn}_{P_{\ell},Q_{\ell}}(\mathbf{Z}_{\ell}),$ (A.11)
+$\mathbf{Z}_{\ell+1}=\mathbf{Z}_{\ell}+\frac{1}{n}\sum_{j=0}^{\ell}\Gamma_{j}^{\ell}\odot\mathbf{R}_{j},$ (A.12)
+where $\Gamma_{j}^{\ell}$ governs the contribution of previous layers, and $\odot$ is the Hadamard (element-wise) product for scaling.
+Our goal is to show that the memory-augmented Transformer with updates given by equations (A.11) and (A.12) can implement $k$ steps of an LFOM, whose general formulation is:
+$\mathbf{w}^{k+1}=\mathbf{w}^{0}+\sum_{i=0}^{k}\Lambda_{i}^{k}\nabla f(\mathbf{w}^{i}),$
+where $\Lambda_{i}^{k}$ are diagonal matrices that scale the gradients $\nabla f(\mathbf{w}^{i})$.
+We will proceed by establishing a correspondence between the variables and updates in the memory-augmented Transformer and those in the LFOM, and by showing that, under appropriate parameter settings, the Transformer updates replicate the LFOM updates.
+The first order of business is to realize that, in the proof of Lemma 1 (A.1), the $\mathbf{w}_{k+1}^{\text{gd}}$ update rule (A.10) is a direct downstream consequence of the $\mathbf{Z}_{\ell+1}$ update rule (2.5), under the parameterization given in equation (2.8).
+Set $\mathbf{R}_{\ell}=\mathrm{Attn}_{P_{\ell},Q_{\ell}}(\mathbf{Z}_{\ell})$ per (A.11).
+Then the consequence of the $\mathbf{Z}_{\ell+1}=\mathbf{Z}_{\ell}+\frac{1}{n}\sum_{j=0}^{\ell}\Gamma_{j}^{\ell}\odot\mathbf{R}_{j}$ update rule is that each $\mathrm{Attn}_{P_{j},Q_{j}}(\mathbf{Z}_{j})$ is coordinate-wise scaled by $\Gamma_{j}^{\ell}\in\mathbb{R}^{(d+1)\times(n+1)}$.
+But if $\mathrm{Attn}_{P_{j},Q_{j}}(\mathbf{Z}_{j})$ is coordinate-wise scaled by $\Gamma_{j}^{\ell}$, then the $\mathbf{Y}_{k+1}$ update rule in (A.1) now instead looks like $\mathbf{Y}_{k+1}=\mathbf{Y}_{k}-\frac{1}{n}\sum_{j=0}^{k}\Gamma_{j}^{k}\big{|}_{d+1}\odot(\mathbf{Y}_{k}M\mathbf{X}_{0}^{\top}A_{k}\mathbf{X}_{0})$, where $\Gamma_{j}^{k}\big{|}_{d+1}$ denotes the $(d+1)$-th row of $\Gamma_{j}^{k}$.
+This is because, by definition, $\mathbf{Y}_{i}$ is the $(d+1)$-th row of $\mathbf{Z}_{i}$ (A.1).
+From the basic $\mathbf{Y}_{k}$ update rule in (A.1), the update formula for $\mathbf{y}_{n+1}^{(k+1)}$ in (A.2) follows as a consequence.
+Except that now, this update formula will include a coordinate-wise scaling as well, which we will denote by $\Lambda_{j}^{k}\in\mathbb{R}^{d}$:
+$\mathbf{y}_{n+1}^{(k+1)}=\mathbf{y}_{n+1}^{(k)}-\frac{1}{n}\sum_{j=0}^{k}\langle(A_{j}\mathbf{\bar{X}}^{\top}\mathbf{\bar{Y}}_{j})\odot\Lambda_{j}^{k},\mathbf{x}_{n+1}\rangle,$
+which in turn leads to $\theta_{k+1}=\theta_{k}-\frac{1}{n}\sum_{j=0}^{k}(A_{j}\mathbf{\bar{X}}\mathbf{\bar{X}}^{\top}(\theta_{j}+\mathbf{w}^{*}))\odot\Lambda_{j}^{k}$ in place of (A.4) and $\mathbf{w}_{k+1}^{\text{gd}}=\mathbf{w}_{k}^{\text{gd}}-\frac{1}{n}\sum_{j=0}^{k}A_{j}\nabla R_{\mathbf{w}^{*}}(\mathbf{w}_{j}^{\text{gd}})\odot\Lambda_{j}^{k}$ in place of (A.5).
+The negative signs can, of course, be incorporated within the $\Lambda_{j}^{k}$s.
+If we simply rewrite $\Lambda_{j}^{k}\in\mathbb{R}^{d}$ as a diagonal matrix in $\mathbb{R}^{d\times d}$, this setup then subsumes the case of diagonal preconditioners $\Lambda_{j}^{k}\in\mathbb{R}^{d\times d}$ acting on the gradients $\nabla R_{\mathbf{w}^{*}}(\mathbf{w}_{j}^{\mathrm{gd}})$, which in the general form looks like:
+$\mathbf{w}_{k+1}^{\mathrm{gd}}=\mathbf{w}_{0}+\sum_{i=0}^{k}\Lambda_{i}^{k}\nabla R_{\mathbf{w}^{*}}(\mathbf{w}_{i}^{\mathrm{gd}}).$
+where $\Lambda_{i}^{k}$ are diagonal matrices.
+Note.
+The memory-augmented Transformer performs exactly these updates in the special case when the preconditioners $A_{j}$ are scalar multiples of the identity.
+If the preconditioners $A_{j}$ are non-trivial, then this architecture performs “LFOM-like" algorithms that lie in a class richer than LFOMs (3.2).
+A.4 Full Proof of Theorem 4.1
+Theorem.
+Consider a multi-head Memformer with $H$ heads, where each head $h$ is parameterized by
+$\bigl{\{}P_{\ell}^{h},\;Q_{\ell}^{h},\;\Gamma_{\ell}^{h}\bigr{\}}_{\ell=1}^{L}.$
+Suppose that in-context prompts $\mathbf{Z}_{0}$ are drawn from a mixture of $M$ Gaussian components,
+$\bigl{\{}\mathcal{N}(\mu_{m},\Sigma_{m})\bigr{\}}_{m=1}^{M}.$
+Assume that, after training, for each mixture component $\mathcal{N}(\mu_{m},\Sigma_{m})$, there exists a head $h_{m}$ whose parameters yield near-optimal first-order performance on data from that component.
+Then, for each component $m$, there exist scalar gating coefficients
+$\{\alpha_{m,h}\}_{h=1}^{H}$
+such that, whenever a new prompt $\mathbf{Z}_{0}$ is drawn from $\mathcal{N}(\mu_{m},\Sigma_{m})$, where $\mathbf{x}^{(i)}\sim\mathcal{N}(\mu_{m},\Sigma_{m})$ i.i.d., and targets $\mathbf{w}^{*}$ from $\mathcal{N}(\mu_{m},\Sigma_{m}^{-1})$, the Memformer can form the linear combination
+$\sum_{h=1}^{H}\alpha_{m,h}\,\mathrm{Head}_{h}(\mathbf{Z}_{0})$
+to achieve near-optimal performance for that component. In particular, if one sets
+$\alpha_{m,h_{m}}=1\quad\text{and}\quad\alpha_{m,h}=0\ \ \text{for all}\ h\neq h_{m},$
+then the multi-head Memformer yields the same near-optimal updates as head $h_{m}$ alone.
+Proof.
+For each head $h\in\{1,\dots,H\}$, let $\mathrm{Head}_{h}(\mathbf{Z}_{0})$ denote the final update (or prediction) produced solely by that head’s parameters
+$\bigl{\{}P_{\ell}^{h},\;Q_{\ell}^{h},\;\Gamma_{\ell}^{h}\bigr{\}}_{\ell=1}^{L}$
+when presented with the in-context prompt $\mathbf{Z}_{0}$.
+By assumption, for each mixture component $m$, there is an associated head $h_{m}$ that is near-optimal for data drawn from $\mathcal{N}(\mu_{m},\Sigma_{m})$.
+We now define gating coefficients
+$\bm{\alpha}_{m}\,=\,(\alpha_{m,1},\,\alpha_{m,2},\,\dots,\,\alpha_{m,H}),$
+to form a linear combination of the $H$ heads’ outputs.
+Specifically, in a forward pass on a new prompt $\mathbf{Z}_{0}$ sampled from $\mathcal{N}(\mu_{m},\Sigma_{m})$, the multi-head Memformer can produce
+$\sum_{h=1}^{H}\alpha_{m,h}\,\mathrm{Head}_{h}(\mathbf{Z}_{0}).$
+Choosing $\alpha_{m,h_{m}}=1$ and $\alpha_{m,h}=0$ for $h\neq h_{m}$ ensures that the combined update coincides exactly with $\mathrm{Head}_{h_{m}}(\mathbf{Z}_{0})$.
+Since $h_{m}$ is by hypothesis near-optimal on that mixture component, the overall performance on $\mathcal{N}(\mu_{m},\Sigma_{m})$ is likewise near-optimal.
+Repeating this argument for each $m$ shows that every mixture component can be matched with the head that specializes in it.
+∎
+A.5 Full Proof of Theorem 5.1
+Theorem.
+Fix a class of LFOMs parameterized by $\theta\in\Theta\subset\mathbb{R}^{p}$, for instance via diagonal preconditioners $\{\Lambda_{i}^{k}\}$.
+Let each LFOM $\theta$ map the in-context prompt $\mathbf{Z}_{0}$ to a final prediction $\hat{y}_{\theta}(\mathbf{Z}_{0})$ for $\mathbf{x}(n+1)$.
+Define the in-context loss
+$\ell(\theta;\mathbf{Z}_{0})\;=\;\Bigl{(}\hat{y}_{\theta}(\mathbf{Z}_{0})-\langle\mathbf{x}(n+1),\mathbf{w}^{*}\rangle\Bigr{)}^{2}.$
+Suppose:
+  1. (i)
+    The domain of $\mathbf{Z}_{0}$, as drawn from distribution $\mathcal{D}$, is such that the squared loss $\ell(\theta;\mathbf{Z}_{0})$ is almost surely bounded by a constant $L>0$ or is sub-Gaussian with scale $\sigma^{2}$.
+  2. (ii)
+    The parameter space $\Theta$ is a compact subset of $\mathbb{R}^{p}$.
+  3. (iii)
+    The mapping $\theta\mapsto\ell(\theta;\mathbf{Z}_{0})$ is continuous (or Lipschitz) for all $\mathbf{Z}_{0}$ in the support of $\mathcal{D}$.
+Draw $N$ i.i.d. prompts $\{\mathbf{Z}_{0}^{(i)}\}_{i=1}^{N}\sim\mathcal{D}$, and let
+$\widehat{\theta}\;=\;\arg\min_{\theta\in\Theta}\;\;\frac{1}{N}\sum_{i=1}^{N}\ell\!\bigl{(}\theta;\,\mathbf{Z}_{0}^{(i)}\bigr{)}.$
+Then, with probability at least $1-\delta$ over the sample draw,
+$\mathbb{E}_{\mathbf{Z}_{0}\sim\mathcal{D}}\bigl{[}\ell(\widehat{\theta};\,\mathbf{Z}_{0})\bigr{]}\;\;\leq\;\;\min_{\theta\in\Theta}\mathbb{E}_{\mathbf{Z}_{0}\sim\mathcal{D}}\bigl{[}\ell(\theta;\,\mathbf{Z}_{0})\bigr{]}\;+\;\epsilon(N,\delta),$
+where $\epsilon(N,\delta)\to 0$ as $N\to\infty$. Thus the learned LFOM parameters $\widehat{\theta}$ generalize well on new prompts from $\mathcal{D}$.
+Proof.
+Define the population risk $R(\theta):=\mathbb{E}_{\mathbf{Z}_{0}\sim\mathcal{D}}[\ell(\theta;\mathbf{Z}_{0})]$ and the empirical risk $\widehat{R}_{N}(\theta):=\frac{1}{N}\sum_{i=1}^{N}\ell(\theta;\mathbf{Z}_{0}^{(i)})$.
+The estimator $\widehat{\theta}$ minimizes $\widehat{R}_{N}(\theta)$ over the compact set $\Theta\subset\mathbb{R}^{p}$.
+Because $p$ is finite and the loss is bounded or sub-Gaussian, classical parametric uniform convergence guarantees apply.
+Concretely, by, for example, the standard Glivenko–Cantelli or Vapnik–Chervonenkis theory for finite-dimensional parameter spaces (see, e.g., Shalev-Shwartz and Ben-David (2014) or Geer (2000)), there is a function $\epsilon(N,\delta)$ going to 00 as $N\to\infty$ such that, with probability at least $1-\delta$, one has
+$\sup_{\theta\in\Theta}\Bigl{|}\,\widehat{R}_{N}(\theta)-R(\theta)\Bigr{|}\;\;\leq\;\epsilon(N,\delta).$
+Since $\widehat{\theta}$ minimizes $\widehat{R}_{N}(\theta)$, a straightforward argument gives
+$R(\widehat{\theta})\;\;=\;\;\widehat{R}_{N}(\widehat{\theta})\;+\;\bigl{(}R(\widehat{\theta})-\widehat{R}_{N}(\widehat{\theta})\bigr{)}\;\;\leq\;\;\widehat{R}_{N}\!\bigl{(}\theta^{*}\bigr{)}\;+\;\epsilon(N,\delta),$
+where $\theta^{*}\in\arg\min_{\theta\in\Theta}R(\theta)$. Another $\epsilon(N,\delta)$-term bounds the difference between $\widehat{R}_{N}(\theta^{*})$ and $R(\theta^{*})$, which yields
+[…]
+Thus, possibly redefining $\epsilon(N,\delta)$ by a small constant factor, we obtain the usual generalization guarantee
+$R(\widehat{\theta})\;\;\leq\;\;\min_{\theta\in\Theta}R(\theta)\;+\;\epsilon(N,\delta).$
+Because $R(\theta)$ is exactly the expected in-context loss in question, this completes the proof. ∎
+Note.
+The boundedness (or sub-Gaussianity) of $\ell(\theta;\mathbf{Z}_{0})$, plus the compactness and continuity assumptions on $\theta$, are key to ensuring uniform convergence in finite-dimensional parameter spaces.
+Such conditions are typically satisfied in standard LFOM settings, e.g.
+when the data are bounded or sub-Gaussian and the square loss is used.
+If data or parameters are unbounded, one can impose appropriate norm constraints and regularity conditions to obtain similar results.
 Appendix B Comparison to Nesterov Accelerated Gradient Method (NAG) and Momentum Gradient Descent (MGD)
@@
 Momentum helps to mitigate oscillations in directions with high curvature, stabilizing the optimization trajectory and leading to faster convergence compared to gradient descent.
-B.3 Memformer LFOMs vs. NAG and MGD
-In our experiments, we observed that memformer LFOMs outperform both NAG and MGD on non-isotropic data.
-Figures 6(a) and 6(b) compare the performance of memformer LFOMs with NAG and MGD, respectively, on the same non-isotropic data.
+B.3 Memformers vs. NAG and MGD
+In our experiments, we observed that Memformers (3.4) outperform both NAG and MGD.
+Figures 14 and 14 compare the performance of Memformer with NAG and MGD, respectively, on the same non-isotropic data.
 As shown, the Memformer achieves faster convergence and much better loss performance compared to both algorithms.
-(a) Nesterov AGM vs. Memformer LFOM on non-isotropic data.
-(b) Momentum GD vs. Memformer LFOM on non-isotropic data.
-Figure 6: Comparison of Nesterov Accelerated Gradient Method (left) and Momentum Gradient Descent (right) vs. Memformer LFOM on non-isotropic data.
+Figure 12: Nesterov AGM vs. LFOM Memformer
+Figure 13: Momentum GD vs. LFOM Memformer
+Figure 14: Comparison of LFOM Memformer with (a) Nesterov AGM and (b) Momentum GD.
+Appendix C LFOM Memformer Experiments With More Than 4 Layers
+In our experiments, we observed that Memformers with more than 4 layers continue to demonstrate impressive performance in learning optimization strategies.
+We conducted experiments with Memformers having up to 7 layers and dimension $d=10$.
+Training beyond this point becomes impractical due to extensive iteration requirements and significant convergence times, which can span several hours.
+This limitation is a consequence of computational constraints (e.g., available GPUs) rather than any inherent deficiency of the Memformer architecture itself.
+Here, $d$ refers to the rank of the square matrix $\mathbf{X}\mathbf{X}^{T}$ in the empirical loss quadratic as described in Equation 2.11.
+1.
+Experiment (Figure 17) (Dimension $d=5$, Layers = 5): As expected, Conjugate Gradient Descent (CGD) converges within $d$ steps due to the dimensionality constraint.
+Remarkably, even though the Memformer only learns general parameters $\mathbf{A}_{\ell}$ (Equation (2.8)) and $\Gamma_{\ell}$ (Equation (3.4)), it manages to keep up with CGD for up to 4 steps, showcasing its efficiency.
+2.
+Experiment (Figure 17) (Dimension $d=10$, Layers = 7): In this case, CGD does not converge until beyond 7 steps, which aligns with theoretical expectations.
+Nevertheless, the Memformer remains highly competitive, matching CGD’s performance for 6 steps and even performing comparably at 7 steps.
+This demonstrates the Memformer’s robust generalization capabilities, even under more complex conditions.
+Figure 15: Memformer performance for $d=5$ with 5 layers.
+Figure 16: Memformer performance for $d=10$ with 7 layers.
+Figure 17: Comparison of Memformer performance for different configurations of depth $d$ and layers.
+Appendix D Experiment on Convergence Verification for Memformer Parameter $\mathbf{A}_{\ell}$ to $\Sigma$
+Our strategy to train the Memformer (3.4) was to first train the $A_{\ell}$’s (2.8) in each layer $\ell$ on the training batch and then to “fine-tune" the $\Gamma_{\ell}$’s on the training batch.
+Therefore, we present here an empirical verification of our results per Theorem 3 in Ahn et al.
+(2024).
+[…]
+$\min_{\{A_{\ell}\}_{\ell=0}^{L-1}}f(A).$
+Let $S\subset\mathbb{R}^{L\times d\times d}$ be defined as follows: $A\in S$ if and only if for all $i=0,\ldots,L-1$, there exist scalars $a_{i}\in\mathbb{R}$ such that $A_{i}=a_{i}\Sigma^{-1}$. Then
+$\inf_{(A,B)\in S}\sum_{i=0}^{L-1}\|\nabla_{A_{i}}f(A,B)\|_{F}^{2}=0,$
+where $\nabla_{A_{i}}f$ denotes the derivative with respect to the Frobenius norm $\|A_{i}\|_{F}$.
+We evaluated the in-context learning (ICL) loss for linear regression with $d=5$ and $n=20$, where $x^{(i)}\sim\mathcal{N}(0,\Sigma)$ and $w_{x}\sim\mathcal{N}(0,\Sigma^{-1})$.
+The covariance $\Sigma$ was generated as $\Sigma=U^{T}DU$, with $U$ being a random orthogonal matrix and $D=\text{diag}(1,1,1/4,1/16,1)$.
+A three-layer linear transformer was trained using ADAM, with $A_{0},A_{1},A_{2}$ initialized as i.i.d.
+Gaussian matrices.
+Each gradient step used minibatches of size 20,000, resampled every 100 steps, and gradients were clipped to 0.01.
+Results were averaged over 5 runs with independent $U$ and $\Sigma$ samples.
+To measure convergence, we computed the normalized Frobenius norm distance:
+$\text{Dist}(M,I):=\min_{\alpha}\frac{\|M-\alpha I\|_{F}}{\|M\|_{F}},$
+where
+$\alpha:=\frac{1}{d}\sum_{i=1}^{d}M[i,i],$
+which quantifies the deviation of $M/\|M\|_{F}$ from a scaled identity.
+The distance $\text{Dist}(\Sigma^{1/2}A_{i}\Sigma^{1/2},I)$, averaged over 5 runs, is shown in Figures 21, 21, and 21 as a function of training iterations.
+Figure 18: $A_{0}$ convergence.
+Figure 19: $A_{1}$ convergence.
+Figure 20: $A_{2}$ convergence.
+Figure 21: Convergence comparison of $A_{0}$, $A_{1}$, and $A_{2}$.